Ű cŚŁäĚŁÏŘ Ç c ŁþŚŁÞŁœ ýùĄŁČŁł ø ñćŚŁśŁŮ ŃŃŃ ŰŚŁäĚŁÏŘ ý c ďøŚŁĂaŁê

ÀŁîeüŁœĄŁþ ğ c ŚŁĺŘ ÂŁŞ üŁĺďŚŁê

üŁþŚŁúŁœ ýĆŁ£ŁÆŁœ

Information Technology – Persian

Information Interchange and Display Mechanism, using Unicode

Final Version Ç c ŁþŚŁÞŁœ ýùĄŁČŁł ø ñćŚŁśŁŮ ŃŃŃ ŰŚŁäĚŁÏŘ ý c ďøŚŁĂaŁê’ ćcďŘÀŁœŚŁţŁĺŘ ü c ŁĂŁê ö cĄŁČŁÆŁČŁÞŁî

"ÀŁîeüŁœĄŁþ ğ c ŚŁĺŘ ÂŁŞ üŁĺďŚŁê Ű cŚŁäĚŁÏŘ

üŁðÀŁĂŁþŚŁÞŁœ ŚŁþ ťŁÞŁĺ

\underline{:ÅŁČŁŐď}
ÓŁþÂŁł ü c ŁţŁãŁĂŁň ù cŚŁÚŁÈŁœŘć  üŁČŁ¡Łþ ,ÇŁŞŚŁŮ

(üŁőŚŁþď ý c ÂŁţŁîć)

\underline{:ŚŁÌŁäŘ}

ÓŁþÂŁł ü c ŁţŁãŁĂŁň ù cŚŁÚŁÈŁœŘć  ćŘÀŁúŁŞ ÀŁČŁĺ ,üŁŞŘÂŁĺ ùćŘěSSŁČŁÆŁ§ÂŁČŁõ ÀŁśŁúŁÔŁĺŘ

(ÂŁŮĄŁČŁşŁõŚŁî ü c ŁĺÀŁĂŁúŁõ Å c ŁœŚŁÆŁČŁó)

ÓŁþÂŁł ü c ŁţŁãŁĂŁň ù cŚŁÚŁÈŁœŘć  ĆŁŞěøď ,ďćŚŁœďĄŁŢ

(ÂŁŮĄŁČŁşŁõŚŁî ü c ŁĺÀŁĂŁúŁõ Å c ŁœŚŁÆŁČŁó)

öÀŁĂŁó ù cŚŁÚŁÈŁœŘć  ÂŁçŁňŘüŁÜŁä ,öŚŁśŁœŚŁą

(üŁőŚŁþď Å c ŁœŚŁÆŁČŁóëĄŁê)

ÓŁþÂŁł ü c ŁţŁãŁĂŁň ù cŚŁÚŁÈŁœŘć  ÀŁČŁõŘ ,üŁœĚŁČŁõ ďŘÀŁÞŁÜŁä

(ÂŁŮĄŁČŁşŁõŚŁî ü c ŁĺÀŁĂŁúŁõ Å c ŁœŚŁÆŁČŁó)

\underline{:ÂŁČŁŞć}

ÓŁþÂŁł ü c ŁţŁãŁĂŁň ù cŚŁÚŁÈŁœŘć  ŘÂŁûě ,üŁûŚŁĂŁŢ

(ÂŁŮĄŁČŁşŁõŚŁî ô cĄŁÜŁä Å c ŁœŚŁÆŁČŁóëĄŁê)

öŘÂŁþŘ ü c ŁţŁãŁĂŁň Ű cŚŁÖŁČŁÖŁ¡ŁŮ ø ćďŘÀŁœŚŁţŁĺŘ ýĆŁÆŁĺĄbŁõ  ÝŁþÂŁõ ,üŁ¡ŁČŁÊŁê

(ÂŁŮĄŁČŁşŁõŚŁî ô cĄŁÜŁä Å c ŁœŚŁÆŁČŁó)

ĆŁ¡ŁÔŁň . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŰŚŻďÀĂõ ťc ĺÂúê

Ť . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ďŚŁţŁÔŁÚŁÈŁČŁŢ

Ÿ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁõÀŁÖŁõ

1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ćÂŁŞďŚŁî ýĆŁĂŁõŘć ø éÀŁû 1

2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁõŘÃŁóŘ â cŁŻŘÂŁõ 2

3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÓŁþďŚŁãŁŮ ø ŰŚŁ§ĚŁÎŁňŘ 3

3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . SSŁţŁõ 1-3

3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÍŁą 2-3

4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁÆŁþĄŁœ 3-3

4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁÆŁþĄŁœĆŁäĄŁÞŁźŁõ 4-3

4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÛŁØŁł 5-3

4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ùćŚŁĺ SS cŁţŁõ 6-3

4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . öćÂŁî ýďŘÁŁðÀŁî 7-3

4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁûćŚŁÞŁœ 4

5 . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ö cĄŁţŁõ ďć ùćŚŁÔŁţŁĺŘ ćcďĄŁõ ýŚŁûĆŁÆŁþĄŁœ 5

6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁóÂŁţŁĂŁî ýŚŁûĆŁÆŁþĄŁœ 1-5

7 . . . . . . . . . . . . . . . . . . . . . . . . . . íÂŁţŁÈŁõ ý c ďŘÁŁðĆŁÎŁÖŁœ Ý cŁŐĚŁä 2-5

9 . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ý c ďŘÁŁðĆŁÎŁÖŁœ Ý cŁŐĚŁä 3-5

9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁőŚŁþď Ý cŁŐĚŁä ø ôŚŁìďŘ 4-5

10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ü c ŁÜŁňŘ éc øÂŁ§ 5-5

13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁäÂŁê éc øÂŁ§ 6-5

14 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ýŚŁûĆŁœŚŁÈŁœ 7-5

15 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . áĄŁĂŁÞŁõ ýŚŁûĆŁÆŁþĄŁœ 8-5

16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ăĄŁÆŁĂŁõ ýŚŁûĆŁÆŁþĄŁœ 9-5

ÓóŘ 17 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁţŁúŁŻøć Ý cŁţŁþďĄŁÚŁóŘ ÓóŘ ťc ĺĄČŢ

18 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ñŚŁÊŁŮŘ Ý cŁţŁþďĄŁÚŁóŘ Š ťc ĺĄČŢ

18 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ñŚŁÊŁŮŘ ýùćď 1-Š

19 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÝŁţŁþďĄŁÚŁóŘ 2-Š

20 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ñŚŁÊŁŮŘ ù cøÂŁð 3-Š

22 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁûďĄŁŮŚŁÚŁČŁó 4-Š

23 . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁûùćŘć ñćŚŁśŁŮ ýŚŁûřŁóŚŁì Ť ťc ĺĄČŢ

24 . . . . . . . . . . . . . . . . . . . . . . . . . . ýÀŁĂŁŞéŘÂŁðŘďŚŁŢ ø ýÀŁĂŁŞÂŁÎŁĺ Ű ťc ĺĄČŢ

25 . . . . . . . . . . . . . . . . . . . . . . . . . . . . ýěďŘÝŁû ø ýěŚŁĺñŚŁõÂŁœ Ÿ ťc ĺĄČŢ

26 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁõŚŁœùęŘø Ž ťc ĺĄČŢ

28 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁûĆŁÆŁþĄŁœ À cŁî Ĳ ťc ĺĄČŢ

31 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁûĆŁÆŁþĄŁœ ô cŚŁœ đ ťc ĺĄČŢ

Š

ďŚŁţŁÔŁÚŁÈŁČŁŢ

ĆŁŁŁŁî "ÀŁŁŁŁîeüŁŁŁŁœĄŁŁŁŁþ ğ c ŚŁŁŁŁĺŘ ÂŁŁŁŁŞ üŁŁŁŁĺďŚŁŁŁŁê Ű cŚŁŁŁŁäĚŁŁŁŁÏŘ Ç c ŁŁŁŁþŚŁŁŁŁÞŁŁŁŁœ ýùĄŁŁŁŁČŁŁŁŁł ø ñćŚŁŁŁŁśŁŁŁŁŮ ŃŃŃ ŰŚŁŁŁŁäĚŁŁŁŁÏŘ ý c ďøŚŁŁŁŁĂaŁŁŁŁê’ ćcďŘÀŁŁŁŁœŚŁŁŁŁţŁŁŁŁĺŘ ø ùÀŁł SSŁþøÀŁŮ ø ĆŁČŁúŁŮ ĆŁÏĄŁŞÂŁõ ýŚŁûöĄŁČŁÆŁČŁÞŁî ďć "ďĄŁÈŁî Œ c ŁČŁŮŚŁõďĄŁÔŁœŘ ü c ŁóŚŁä ý c ŘďĄŁł’ Íc ŁĺĄŁŮ öŊ Å c ŁþĄŁĂŁÈŁČŁŢ ÀŁŁČŁŁþŚbŁŁŮ ćcďĄŁŁõ 1381/2/18 ă c ďĄŁŁõ ŚŁŁûùćŘć ý c ďøŊÂŁŁê ø ĆŁŁœŚŁŁþŘď ćcďŘÀŁŁœŚŁŁţŁŁĺŘ ü c ŁŁÜŁŁõ ýĆŁŁţŁŁČŁŁÞŁŁî ýĆŁŁÆŁŁÜŁŁŻ SSŁŁČŁŁÞŁŁźŁŁĂŁŁŢ ďć ø ćďŘÀŁœŚŁţŁĺŘ ýĆŁÆŁĺĄbŁõ Ű cŘďÂŁÖŁõ ø SSŁČŁœŘĄŁì đ c ĚŁňŘ ö cĄŁœŚŁì 3 ýùćŚŁõ Œ c Łþ À cŁĂŁŞ ćcŚŁĂŁţŁĺŘ ĆŁŞ ŒŁĂŁþŘ ,ťŁĺŘ ĆŁţŁêÂŁð ďŘÂŁì

.ćĄŁłüŁõ ÂŁÈŁţŁĂŁõ öŘÂŁþŘ ü c ŁÜŁõ ćcďŘÀŁœŚŁţŁĺŘ ö cŘĄŁĂŁäĆŁŞ 1371 ù cŚŁõSSŁÞŁúŁŞ Š c ĄŁÊŁõ ,öŘÂŁþŘ ü c ŁţŁãŁĂŁň Ű cŚŁÖŁČŁÖŁ¡ŁŮ

ø ,ôĄŁÜŁä ,âŁþŚŁĂŁň ýĆŁĂŁČŁõě ďć üŁœŚŁúŁŻ ø üŁÜŁõ ýŚŁûťŁêÂŁÈŁČŁŢ ø ŰĎĄŁ¡ŁŮ ŚŁŞ üŁÚŁĂŁûŚŁÞŁû ø üŁõŚŁÚŁÞŁû Ð c ŁÔŁ§ ýŘÂŁŞ ýŘÂŁŁŞ ĆŁŁî ýćŚŁŁúŁŁĂŁŁÈŁŁČŁŁŢ ĆŁŁœĄŁŁðÂŁŁû ø ÀŁŁł ÀŁŁĂŁŁûŘĄŁŁą ÂŁŁÑŁŁœ À cŁŁþÀŁŁźŁŁŮ ôøÃŁŁó â cŁŁìŘĄŁŁõ ďć öŘÂŁŁþŘ ü c ŁŁÜŁŁõ ýŚŁŁûćďŘÀŁŁœŚŁŁţŁŁĺŘ ,ŰŚŁŁõÀŁŁą ĆŁŻĄŁŮ ćcďĄŁõ ĆŁÏĄŁŞÂŁõ ü c ŁĂŁê ýŚŁûöĄŁČŁÆŁČŁÞŁî ďć ÂŁÑŁœ À cŁþÀŁźŁŮ ô cŚŁÚŁĂŁû ďć ,ćĄŁł ĆŁŐŘďŘ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ Û cŁČŁÞŁØŁŮ ŚŁþ đĚŁňŘ ŚŁûöŊ Â cŁÑŁœÀ cŁþÀŁźŁŮ SSŁþÂŁąŊ ěŘ ùďŘĄŁÞŁû ÀŁþŚŁŞ öŘÂŁþŘ ýŚŁûćďŘÀŁœŚŁţŁĺŘ ĆŁŞ ĆŁãŁŻŘÂŁõ ýŘÂŁŞ SSŁþŘÂŁŞŚŁĂŁŞ .ťŁêÂŁð ÀŁûŘĄŁą ďŘÂŁì

.ćÂŁî ùćŚŁÔŁţŁĺŘ

ďć ,ĆŁŁãŁŁõŚŁŁŻ ýŚŁŁûěŚŁŁČŁŁœ ø ćĄŁŁŻĄŁŁõ Íc ŁŁþŘÂŁŁł ĆŁŁŞ ĆŁŁŻĄŁŁŮ SS cŁŁÞŁŁő ĆŁŁî ťŁŁĺŘ ùÀŁŁł üŁŁãŁŁĺ ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ SS cŁŁþøÀŁŁŮ ø ĆŁŁČŁŁúŁŁŮ ďć ø üŁŁţŁŁãŁŁĂŁŁň ýŚŁŁûďĄŁŁÈŁŁî ü c ŁŁÜŁŁõ ýŚŁŁûćďŘÀŁŁœŚŁŁţŁŁĺŘ ø üŁŁÜŁŁÜŁŁÞŁŁóŘSSŁŁČŁŁŞ ýŚŁŁûćďŘÀŁŁœŚŁŁţŁŁĺŘ ø ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ SS cŁŁČŁŁŞ öŚŁŁØŁŁõŘ À cŁŁ§

.ćĄŁł ćŚŁźŁþŘ üŁÚŁĂŁûŚŁÞŁû ĆŁţŁêÂŁÈŁČŁŢ

:ťŁĺŘ ÂŁþě đ c ÂŁł ĆŁŞ ĆŁţŁêď ďŚŁî ĆŁŞ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ýĆŁČŁúŁŮ ýŘÂŁŞ ĆŁî üŁãŁŻŘÂŁõ ø âŁŞŚŁĂŁõ

1. The Unicode Consortium, The Unicode Standard, Version 3.2.0, defined by: The Unicode Standard, Version 3.0, Addison-Wesley, 2000, as amended by the Unicode Standard Annex #27: Unicode 3.1

(http://www.unicode.org/unicode/reports/tr27/)

and by the Unicode Standard Annex #28: Unicode 3.2

(http://www.unicode.org/unicode/reports/tr28/).

2. ISO 10646-1:2000 Information Technology – Universal Multiple-Octet Coded Character Set (UCS) – Part 1: Architecture and Basic Multilingual Plane.

3. Dave Ragget, Arnaud Le Hors, Ian Jacobs, “HTML 4.01 Specification”, World Wide

Web Consortium, December 1999.

4. Tim Bray, Jean Paoli, C. M. Sperberg-McQueen, Eve Maler, “Extensible Markup Language (XML) 1.0”, World Wide Web Consortium, Second Edition, October 2000.

5. Martin J. Dürst, François Yergeau, Richard Ishida, Misha Wolf, Asmus Freytag, Tex Texin, “Character Model for the World Wide Web 1.0”, World Wide Web Consortium, Working Draft, April 2002.

.üŁĺďŚŁê ü c ŁţŁČŁŞ 8 Ű cŚŁäĚŁÏŘ ñ cćŚŁśŁŮ À cŁîe 1372 ñ cŚŁĺ :3342 öŘÂŁþŘ ü c ŁÜŁõ ćcďŘÀŁœŚŁţŁĺŘ .6 .üŁĺďŚŁê ö cŚŁŞě ĆŁŞ Ű cŚŁäĚŁÏŘ ñ cćŚŁśŁŮ À cŁîe 1367 ñ cŚŁĺ :2900 öŘÂŁþŘ ü c ŁÜŁõ ćcďŘÀŁœŚŁţŁĺŘ .7 .ÂŁþÂŁ¡ŁŮ ýŚŁûSSŁČŁłŚŁõ ďć üŁĺďŚŁê éc øÂŁ§ 1351 ñ cŚŁĺ :820 öŘÂŁþŘ ü c ŁÜŁõ ćcďŘÀŁœŚŁţŁĺŘ .8

.1378 ,üŁĺďŚŁê Š c ćŘ ø öŚŁŞě ö cŚŁţŁÆŁÚŁĂŁûÂŁê ,üŁĺďŚŁê Íc Łą ď cĄŁţŁĺć .9

.1372 ,ôøć Ç c ŁþŘÂŁþø ,üŁûŚŁÚŁÈŁœŘć Â cŁÈŁœ Ã cŁîÂŁõ ,ĆŁõŚŁœùĄŁČŁł .10

,ÓŁŁþÂŁŁł ü c ŁŁţŁŁãŁŁĂŁŁň ù cŚŁŁÚŁŁÈŁŁœŘć ,ŰŚŁŁśŁŁĺŚŁŁ¡ŁŁõ Ã cŁŁîÂŁŁõ ,"ĆŁŁØŁŁśŁŁł ďć üŁŁĺďŚŁŁê’ ù cøÂŁŁð ü c ŁŁŮŚŁŁÖŁŁČŁŁÖŁŁ¡ŁŁŮ ýŚŁŁûùęøÂŁŁŢ ÿ c ŁŁþŚŁŁţŁŁœ .11 .1381 ŚŁŮ 1377 ,öŘÂŁúŁŮ

Ű

ĆŁõÀŁÖŁõ

ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ŚŁŞ üŁþŚŁĂŁłŊ

ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ .ťŁŁĺŘ öĄŁŁţŁŁõ ø ŚŁŁûĆŁŁÆŁŁþĄŁŁœ ý c ďŘÁŁŁðÀŁŁîe ýŘÂŁŁŞ üŁŁœŚŁŁúŁŁŻ ýŘùĄŁŁČŁŁł (Unicode) ÀŁŁîeüŁŁœĄŁŁþ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ đ c ĄŁŁŁÎŁŁŁĺ ďć Řď ŰŚŁŁŁäĚŁŁŁÏŘ ñ cćŚŁŁŁśŁŁŁŮ ĆŁŁŁî ÀŁŁŁĂŁŁŁîüŁŁŁõ ÉŁŁŁ£ŁŁŁÈŁŁŁõ ĆŁŁŁœŚŁŁŁŞěÀŁŁŁĂŁŁŁİ ö cĄŁŁŁţŁŁŁõ ý c ďŘÁŁŁŁðÀŁŁŁîe ýŘÂŁŁŁŞ ÙŁŁŁĂŁŁŁûŚŁŁŁÞŁŁŁû üŁŁŁłøď ø HTML Û cŁŁČŁŁśŁŁì ěŘ ,ťŁŁœÂŁŁţŁŁĂŁŁþŘ ýŚŁŁûćďŘÀŁŁœŚŁŁţŁŁĺŘ ŋ c ÂŁŁêÇŁŁČŁŁŢ ý c ďŘÁŁŁðÀŁŁîe ÀŁŁîüŁŁœĄŁŁþ .ćěŚŁŁĺüŁŁõ ÂŁŁÆŁŁČŁŁõ üŁŁÜŁŁÜŁŁÞŁŁóŘSSŁŁČŁŁŞ Ű cŚŁŁŁśŁŁŁŹ .ćĄŁŁŁłüŁŁŁõ üŁŁŁœŚŁŁŁśŁŁŁČŁŁŁţŁŁŁÈŁŁŁŢ ýěøÂŁŁŁõŘ ý c ěŚŁŁŁĺĆŁŁŁõŚŁŁŁœÂŁŁŁŞ ýŚŁŁŁûöŚŁŁŁŞě ø ŚŁŁŁûÛŁŁŁõŚŁŁŁäÝŁŁŁţŁŁŁÆŁŁŁČŁŁŁĺ ýĆŁŁŁČŁŁŁÜŁŁŁî ďć ø ťŁŁŁĺŘ XML ýĆŁŁÜŁŁÞŁŁŻ ěŘ ,ÀŁŁČŁŁóĄŁŁŮ ýŚŁŁûĆŁŁĂŁŁþÃŁŁû ö cÀŁŁł ÝŁŁî ø ŚŁŁûďŘÃŁŁêŘôÂŁŁœ ö cÀŁŁłùćŚŁŁĺ ,öĄŁŁţŁŁõ ü c ŁŁÜŁŁÜŁŁÞŁŁóŘSSŁŁČŁŁŞ ñ cćŚŁŁśŁŁŮ ö cŚŁŁØŁŁõŘ ,ŚŁŁûùćŘć

.ťŁĺŘ ŰŚŁäĚŁÏŘ ý c ďøŊSSŁê ťc ŁãŁĂŁň ýŘÂŁŞ ÀŁîüŁœĄŁþ ýŚŁþŘÃŁõ

,ĆŁÆŁþĄŁœ öĄŁČŁÜŁČŁõ ŒŁþ ěŘ ÇŁČŁŞ ťc ŁČŁêÂŁÒ ŚŁŞ ø ĆŁţŁêď ÂŁŮŘÂŁê ďŚŁČŁÆŁŞ üŁţŁČŁŞ 8 ćcøÀŁ¡Łõ ýŚŁûĆŁÆŁþĄŁœĆŁäĄŁÞŁźŁõ ěŘ ÀŁîüŁœĄŁþ ö cŚŁŁŞě ø ÍŁŁą Š c ŚŁŁ£ŁŁţŁŁœŘ ýŘÂŁŁŞ ,ùøĚŁŁäĆŁŁŞ .ÀŁŁĂŁŁîüŁŁõ ÝŁŁûŘÂŁŁê Řď ŚŁŁČŁŁœć ý c ďŚŁŁţŁŁłĄŁŁœ ýŚŁŁûöŚŁŁŞě ýĆŁŁČŁŁÜŁŁî ý c ďŘÁŁŁðÀŁŁîe ö cŚŁŁØŁŁõŘ ýŚŁûĆŁÆŁþĄŁœ ,üŁþŚŁśŁÔŁóŘ ýŚŁûĆŁÆŁþĄŁœ ŚŁŞ üŁœŚŁÆŁØŁþ ď cŚŁţŁêď ÀŁîüŁœĄŁþ .ćďŘÀŁœ üŁóÂŁţŁĂŁî ýŚŁûÀŁîe ěŘ ùćŚŁÔŁţŁĺŘ ĆŁŞ ýěŚŁČŁœ ,SSŁţŁõ .ÀŁĂŁîüŁõ ÝŁûŘÂŁê ÂŁÚŁþÀŁØŁþ ŚŁŞ ŕc ĚŁţŁąŘ ďć Řď ŚŁûöŊ ěŘ ùćŚŁÔŁţŁĺŘ ö cŚŁØŁõŘ ĆŁî ,ćďŘć ŚŁûĆŁœŚŁÈŁœ ø ŚŁûćŚŁÞŁœ ø ,ďŚŁÚŁœĆŁÈŁþÀŁœŘ ,ťŁĺŘ ùćĄŁŞ ñĄŁÞŁãŁõ ĆŁŞŚŁÈŁõ ýŚŁûćďŘÀŁœŚŁţŁĺŘ ďć ĆŁî ĆŁÆŁþĄŁœ ÂŁû ýŘÂŁŞ ôŚŁœ ø ýćÀŁä À cŁîe SS cŁČŁČŁãŁŮ ÂŁŞ ùøĚŁä ,ÀŁîüŁœĄŁþ ĆŁŁŁÜŁŁŁÞŁŁŁŻ öŊ ěŘ ĆŁŁŁî ,ÀŁŁŁĂŁŁŁîüŁŁŁõ SSŁŁŁČŁŁŁõŚbŁŁŁŮ ťŁŁŁĺŘ ôěĎ öĄŁŁŁţŁŁŁõ Ç c ŁŁŁþŚŁŁŁÞŁŁŁœ ø ľěŘćÂŁŁŁŢ ýŘÂŁŁŁŞ ĆŁŁŁî ÃŁŁŁČŁŁŁœ Řď ýÂŁŁŁţŁŁŁÈŁŁŁČŁŁŁŞ Ű cŚŁŁŁäĚŁŁŁÏŘ

.ćÂŁî ùďŚŁłŘ üŁþŚŁśŁÔŁóŘ ýŚŁûüŁðÄŁþø ø ĆŁÆŁþĄŁœ ťc ŁúŁŻ ĆŁŞ öŘĄŁŮüŁõ

ćcĄŁŁŻĄŁŁõ ýŚŁŁûÂŁŁţŁŁÆŁŁŞ ýŘÂŁŁŞ UTF-8 :ÀŁŁĂŁŁîüŁŁõ ÝŁŁûŘÂŁŁê ŰŚŁŁäĚŁŁÏŘ ý c ěŚŁŁĺùÂŁŁČŁŁąč ø ñćŚŁŁśŁŁŮ ýŘÂŁŁŞ řŁŁóŚŁŁì ĆŁŁĺ ÀŁŁîeüŁŁœĄŁŁþ ýŘÂŁŁŁŞ UTF-16 ,(ťŁŁŁœÂŁŁŁţŁŁŁĂŁŁŁþŘ ĆŁŁŁÜŁŁŁÞŁŁŁŻ ěŘ ,ASCII ćcďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ ÂŁŁŁŞ üŁŁŁĂŁŁŁţŁŁŁśŁŁŁõ ýŚŁŁŁûÍŁŁŁČŁŁŁ¡ŁŁŁõ ýŘÂŁŁŁŞ řŁŁŁĺŚŁŁŁĂŁŁŁõ) üŁŁŁţŁŁŁČŁŁŁŞ 8 ďć ÀŁŁŁîüŁŁŁœĄŁŁŁþ ćcďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ ,SSŁŁŁþŘ ÂŁŁŁŞ ùøĚŁŁŁä .üŁŁŁţŁŁŁČŁŁŁŞ 32 ýŚŁŁŁûÍŁŁŁČŁŁŁ¡ŁŁŁõ ýŘÂŁŁŁŞ UTF-32 ø ,üŁŁŁţŁŁŁČŁŁŁŞ 16 ýŚŁŁŁûÍŁŁŁČŁŁŁ¡ŁŁŁõ ďć .ťŁĺŘ ñćŚŁãŁõ ø ÙŁĂŁûŚŁÞŁû ISO/IEC 10646 ü c ŁÜŁÜŁÞŁóŘSSŁČŁŞ ćcďŘÀŁœŚŁţŁĺŘ ŚŁŞ fĚŁõŚŁî ŚŁûĆŁÆŁþĄŁœ ĆŁŞ ÀŁîe É c ŁČŁÊŁ£ŁŮ ďŚŁðěŚŁĺ fĚŁõŚŁî ÃŁČŁœ ISO/IEC 10646 ćcďŘÀŁœŚŁţŁĺŘ ŚŁŞ ,ÀŁĂŁî ýøďüŁŢ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ěŘ ĆŁî ýćÂŁŞďŚŁî ÂŁû ,âŁìŘø

.ťŁĺŘ

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ 1 Û cŁÊŁê ĆŁŞ ,ÂŁţŁÈŁČŁŞ Ű cŚŁäĚŁÏŘ ýŘÂŁŞ

Ÿ

ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ü c ŁõĄŁÞŁä ď cŚŁţŁąŚŁĺ

ùćøÀŁŁ¡ŁŁõ SSŁŁþŘ .ÀŁŁûćüŁŁõ ńŚŁŁÊŁŁţŁŁąŘ 1’114’111 ŚŁŁŮ 0 ěŘ üŁŁþŚŁŁţŁŁØŁŁþ ćcÀŁŁä ĆŁŁÆŁŁþĄŁŁœ ÂŁŁû ĆŁŁŞ ,ÀŁŁîüŁŁœĄŁŁþ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ Â cŁŁŁůŁŁŁîŘ ø ćďŘć ôŚŁŁŁœ ĆŁŁŁþŚŁŁŁŢ ýĆŁŁŁ¡ŁŁŁÔŁŁŁň ñøŘ ýĆŁŁŁ¡ŁŁŁÔŁŁŁň .ćĄŁŁŁłüŁŁŁõ ÝŁŁŁČŁŁŁÆŁŁŁÖŁŁŁŮ ýŘĆŁŁŁÆŁŁŁþĄŁŁŁœ 65’536 ýĆŁŁŁ¡ŁŁŁÔŁŁŁň 17 ĆŁŁŁŞ

.ćÂŁČŁðüŁõ ÂŁŞ ďć Řď ŚŁČŁœć ýùÀŁœě ýŚŁûöŚŁŞě ďć ùćŚŁÔŁţŁĺŘ ćcďĄŁõ ýŚŁûĆŁÆŁþĄŁœ

ýŘÂŁŁŞ ÀŁŁœŘĄŁŁŮüŁŁõ ĆŁŁî ÀŁŁĂŁŁîüŁŁõ ÉŁŁ£ŁŁÈŁŁõ üŁŁňĄŁŁÊŁŁą ýùćŚŁŁÔŁŁţŁŁĺŘ ýŘÂŁŁŞ ÃŁŁČŁŁœ Řď ĆŁŁÆŁŁþĄŁŁœ ďŘÃŁŁûÀŁŁň ěŘ ÇŁŁČŁŁŞ ÀŁŁîüŁŁœĄŁŁþ

.ćøď ďŚŁîĆŁŞ ŰŚŁäĚŁÏŘ ñ cćŚŁśŁŮ ýŘÂŁŞ SSŁČŁêÂŁÏ Õ cŁêŘĄŁŮ ŚŁŞ ŚŁþ ,üŁÜŁąŘć ý c ěŚŁĺùÂŁČŁąč

.ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ Řď ĆŁÆŁþĄŁœ Û cŁØŁł ŚŁþ ŚŁĂŁãŁõ ĆŁî ÀŁûćüŁõ ÉŁČŁÊŁ£ŁŮ üŁÊŁ£ŁÈŁõ ø ŚŁţŁØŁþ ô cŚŁœ ĆŁÆŁþĄŁœ ÂŁû ĆŁŞ ÀŁîüŁœĄŁþ Řď ĆŁÆŁþĄŁœ öŊ ýŚŁĂŁãŁõ ĆŁî ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ Řď ýŘüŁŮŚŁäĚŁÏŘ ŚŁþ üŁõŘÃŁóŘ ýŚŁûüŁðÄŁþø ĆŁÆŁþĄŁœ ÂŁû ýŘÂŁŞ ,ùøĚŁäĆŁŞ

.ÀŁĂŁĂŁîüŁõ SSŁČŁãŁõ

ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ýĆŁČŁúŁŮ ýùĄŁČŁł

À cŁŁŁŁÊŁŁŁŁì ø ťŁŁŁŁĺŘ ùÀŁŁŁŁł ĆŁŁŁŁČŁŁŁŁúŁŁŁŁŮ (üŁŁŁŁĺďŚŁŁŁŁê ö cŚŁŁŁŁŞě ý c ďŚŁŁŁŁţŁŁŁŁłĄŁŁŁŁœ Ű cďĄŁŁŁŁň) üŁŁŁŁĺďŚŁŁŁŁê Íc ŁŁŁŁą ñ cćŚŁŁŁŁśŁŁŁŁŮ ýŘÂŁŁŁŁŞ ÂŁŁŁŁőŚŁŁŁŁ§ ćcďŘÀŁŁŁŁœŚŁŁŁŁţŁŁŁŁĺŘ ĆŁØŁÜŁŞ .ćďŘÀŁœ Řď ÍŁą SSŁþŘ ýŚŁûĆŁÆŁþĄŁœ ýùÂŁþŘć ö cćÂŁî ćøÀŁ¡Łõ ŚŁþ ,ľďŚŁÚŁœ ýŘÂŁŞ üŁãŁŻÂŁõ ýùĄŁČŁł ö cćÂŁî ÉŁ£ŁÈŁõ ýŚŁûñĄŁìÛ cŁÖŁœ Û cŁõŚŁł ö cĄŁţŁõ ŚŁþ üŁĺďŚŁê ö cĄŁţŁõ ďć ùÂŁõěøď ýùćŚŁÔŁţŁĺŘ ćcďĄŁõ ýŚŁûĆŁÆŁþĄŁœ ýĆŁČŁÜŁî ťŁĺŘ ùÀŁł ľĚŁŮ

.ÀŁœĄŁł ÂŁîč ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ďć ,ÀŁĂŁłŚŁŞ ćĄŁŻĄŁõ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ďć ĆŁî üŁŮďĄŁň ďć ,üŁśŁûÁŁõ

ŚŁŁŞ ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ Û cŁŁõŚŁŁî ý c ďŚŁŁðěŚŁŁĺ ěŘ öŚŁŁĂŁŁČŁŁÞŁŁÏŘ ýŘÂŁŁŞ Řď SSŁŁØŁŁÞŁŁõ ľc ĚŁŁŮ ťc ŁŁþŚŁŁúŁŁœ ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ ü c ŁŁĂŁŁê ýĆŁŁţŁŁČŁŁÞŁŁî

.ťŁĺŘ ùćŘć ôŚŁźŁœŘ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ

,ÂŁŁŁőŚŁŁŁ§ ćcďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ řc ŁŁŁþĄŁŁŁÊŁŁŁŮ ýŘÂŁŁŁŞ ćďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ ü c ŁŁŁÜŁŁŁõ ýĆŁŁŁţŁŁŁČŁŁŁÞŁŁŁî ýĆŁŁŁÆŁŁŁÜŁŁŁŻ SSŁŁŁČŁŁŁóøŘ ěŘ ÅŁŁŁŢ ŃŃŃ ýďøŊćŚŁŁŁþ ÂŁŁČŁŁąŘ ýĆŁŁ£ŁŁÆŁŁœ ŚŁŁŞ ÂŁŁőŚŁŁ§ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ .ÀŁŁł ÂŁŁÈŁŁţŁŁĂŁŁõ 1381/1/7 ¿ c ŁŁþďŚŁŁŮ ďć ÀŁŁîüŁŁœĄŁŁþ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ 3.2 ýĆŁŁ£ŁŁÆŁŁœ ,ďĄŁÈŁî Œ c ŁČŁŮŚŁõďĄŁÔŁœŘ ü c ŁóŚŁä ýŘďĄŁł ťc ŁĺŘĄŁąďć ĆŁŞ ,ÂŁČŁąŘ ýĆŁ£ŁÆŁœ .ťŁĺŘ ďŚŁðěŚŁĺ fĚŁõŚŁî ÃŁČŁœ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ěŘ ùćŚŁŁÔŁŁţŁŁĺŘ ĳ c ŁŁČŁŁ¡ŁŁň ýùĄŁŁČŁŁł .ťŁŁĺŘ ĆŁŁţŁŁêÂŁŁð ÂŁŁÑŁŁœ ďć ,U+FDFC À cŁŁî ŚŁŁŞ ,"ñŚŁŁþď’ ťc ŁŁõĚŁŁä ýŘÂŁŁŞ ÃŁŁČŁŁœ ýŘĆŁŁÆŁŁþĄŁŁœ

.ÀŁł ÀŁûŘĄŁą ÂŁÈŁţŁĂŁõ ÂŁőŚŁ§ ćcďŘÀŁœŚŁţŁĺŘ ÂŁŞ ýŘĆŁÞŁČŁÞŁő ďć ĆŁÆŁþĄŁœ SSŁþŘ

Ž

üŁĺďŚŁê Ű cŚŁäĚŁÏŘ Ç c ŁþŚŁÞŁœ ýùĄŁČŁł ø ñćŚŁśŁŮ ŃŃŃ ŰŚŁäĚŁÏŘ ý c ďøŚŁĂaŁê

ÀŁîeüŁœĄŁþ ğ c ŚŁĺŘ ÂŁŞ

ćÂŁŞďŚŁî ýĆŁĂŁõŘć ø éÀŁû 1

ø ÀŁŁŁŁŁŁîeüŁŁŁŁŁŁœĄŁŁŁŁŁŁþ ćcďŘÀŁŁŁŁŁŁœŚŁŁŁŁŁŁţŁŁŁŁŁŁĺŘ øć ěŘ ĳŁŁŁŁŁŁČŁŁŁŁŁŁ¡ŁŁŁŁŁŁň ýùćŚŁŁŁŁŁŁÔŁŁŁŁŁŁţŁŁŁŁŁŁĺŘ ýùĄŁŁŁŁŁŁČŁŁŁŁŁŁł SS cŁŁŁŁŁŁČŁŁŁŁŁŁČŁŁŁŁŁŁãŁŁŁŁŁŁŮ ćďŘÀŁŁŁŁŁŁœŚŁŁŁŁŁŁţŁŁŁŁŁŁĺŘ SSŁŁŁŁŁŁþŘ SS cŁŁŁŁŁŁþøÀŁŁŁŁŁŁŮ ěŘ éÀŁŁŁŁŁŁû ýŚŁŁŁûÝŁŁŁţŁŁŁÆŁŁŁČŁŁŁĺ ďć ĆŁŁŁî ťŁŁŁĺŘ ĆŁŁŁœŚŁŁŁŞěÀŁŁŁĂŁŁŁİ ö cĄŁŁŁţŁŁŁõ ü c ŁŁŁĺďŚŁŁŁê ýŚŁŁŁûťŁŁŁÞŁŁŁÆŁŁŁì ø üŁŁŁĺďŚŁŁŁê ö cĄŁŁŁţŁŁŁõ ýŘÂŁŁŁŞ ISO/IEC 10646 ,ÀŁłŚŁŞ ĆŁţŁłŘć Řď ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ďć ěŚŁČŁœ ćcďĄŁõ ü c ŁłěŘćÂŁŢ ýŚŁûüŁþŚŁœŘĄŁŮ ĆŁî ýÂŁÚŁþć Ý cŁţŁÆŁČŁĺ ĆŁœĄŁðÂŁû ŚŁþ ,ýÂŁŮĄŁČŁşŁõŚŁî

.ćøďüŁõ ďŚŁîĆŁŞ

ö cŚŁŁŞě ý c ďŚŁŁţŁŁłĄŁŁœ Ű cďĄŁŁň ýĆŁŁŐŘďŘ ø ,ćøďø ,ýěŚŁŁĺùÂŁŁČŁŁąč ,ľěŘćÂŁŁŢ ,ñćŚŁŁśŁŁŮ ,ñŚŁŁÖŁŁţŁŁœŘ ,ÇŁŁþŚŁŁÞŁŁœ ďć ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ

.ćøďüŁõ ďŚŁî ĆŁŞ öŊ ýŘÂŁŞ ôěĎ ýŚŁûćŚŁÞŁœ ø üŁĺďŚŁê

ÀŁîüŁœĄŁþ ô cŚŁÚŁÞŁû ýŚŁûćďŘÀŁœŚŁţŁĺŘ ğ c ŚŁĺŘ ÂŁŞ Řď üŁĺďŚŁê Ű cŚŁäĚŁÏŘ Ç c ŁþŚŁÞŁœ ø ñćŚŁśŁŮ ĳ c ŁČŁ¡Łň ýùĄŁČŁł ,ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ

.ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ ISO/IEC 10646 ø

:ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ

,ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ Řď üŁĺďŚŁê Íc Łą ďć ùćŚŁÔŁţŁĺŘ ćcďĄŁõ ýŚŁûĆŁÆŁþĄŁœ ŚŁŞ ÂŁÒŚŁĂŁţŁõ À cŁîeø üŁĂŁãŁõ ,ôŚŁœ •

,ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ Řď üŁĺďŚŁê éc øÂŁ§ ñ cŚŁÊŁŮŘ ýùĄŁČŁł ø ,ĆŁţŁúŁŻøć ö cĄŁţŁõ ďć Řď ŚŁûĆŁÆŁþĄŁœ Ç c ŁþŚŁÞŁœ ýùĄŁČŁł •

,ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ ŚŁûùćŘć ñ cćŚŁśŁŮ ýŘÂŁŞ Řď ISO/IEC 10646 ø ÀŁîüŁœĄŁþ Ó c ŁÜŁţŁ£Łõ ýŚŁûřŁóŚŁì •

,ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ Řď ŚŁûÀŁĂŁŞ ø ŚŁûÂŁÎŁĺ ýŚŁúŁţŁœŘ ö cćÂŁî SSŁČŁãŁõ ýùĄŁČŁł •

.ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ ýěďŘÝŁû Â cŁÑŁœ ěŘ Řď ýŘĆŁÆŁþĄŁœ ýŚŁûĆŁţŁłď ýĆŁÆŁþŚŁÖŁõ ýùĄŁČŁł •

ćěŘćÂŁŢüŁÞŁœ ŚŁûöŊ ĆŁŞ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ĆŁî üŁÜŁŐŚŁÆŁõ ěŘ üŁÌŁãŁŞ

:ćěŘćÂŁŢüŁÞŁœ ÂŁþě ćcďŘĄŁõ ĆŁŞ ÂŁőŚŁ§ ćcďŘÀŁœŚŁţŁĺŘ ŚŁûùćŘć ý c ěŚŁĺćďŘø ýŚŁûùĄŁČŁł •

ĆŁœŚŁŞěÀŁĂŁİ ø üŁĺďŚŁê Ű cŘďŚŁśŁä ýěŚŁĺřŁŮÂŁõ •

öĄŁţŁõ ýÀŁĂŁŞÂŁÎŁĺ ø üŁĂŁØŁłÂŁÎŁĺ ýùĄŁČŁł •

ŚŁûSSŁţŁõÂŁþě éÁŁ§ ø ,Žďć ,ÇŁþŘÂŁþø ýùĄŁČŁł •

ÝŁźŁ§ÝŁî Ű cďĄŁňĆŁŞ ŚŁûöŊ ýĆŁóćŚŁśŁõ ŚŁþ ,öĄŁţŁõ ý c ěŚŁĺùćÂŁÈŁê •

ŚŁûSSŁţŁõÂŁþě ø öĄŁţŁõ ö cŚŁŞě öćÂŁî ÉŁ£ŁÈŁõ •

,ÀŁĂŁîüŁõ SSŁČŁãŁõ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ĆŁî ĆŁİöŊ ŚŁŮ ťŁĺŘ ĆŁţŁêÂŁð ŰďĄŁň SSŁØŁÞŁõ ľc ĚŁŮ ťc ŁþŚŁúŁœ ŃŃŃ 1 ýďøŊćŚŁþ ťc ŁÞŁÆŁì ěŘ ýøďüŁŢ ćĄŁł ťŁŞŚŁŹ ĆŁî üŁŮďĄŁň ďć .ÀŁłŚŁśŁœ ISO/IEC 10646 ø ÀŁîüŁœĄŁþ ýŚŁûćďŘÀŁœŚŁţŁĺŘ éc ĚŁąÂŁŞ ćcďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ øć Â cŁŁŁČŁŁŁČŁŁŁçŁŁŁŮ ŚŁŁŁþ ,ÀŁŁŁĂŁŁŁîüŁŁŁõ ďŚŁŁŁðěŚŁŁŁĺŚŁŁŁœ ćďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ øć öŊ ŚŁŁŁŞ Řď ŚŁŁŁûćÂŁŁŁŞďŚŁŁŁî ,ćďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ SSŁŁŁþŘ ěŘ üŁŁŁÊŁŁŁ£ŁŁŁÈŁŁŁõ ø ùćĄŁŞ ÛŁÏŚŁŞ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ěŘ (ťŁÞŁÆŁì öŊ ÍŁÖŁê ø) ťŁÞŁÆŁì öŊ ,ćĄŁł ýďŚŁðěŚŁĺŚŁœ ű c ŁäŚŁŞ ùÀŁĂŁþŊ ďć ÂŁîÁŁóŘëĄŁê ,üŁŁŮďĄŁŁň SSŁŁČŁŁĂŁŁİ ďć .ćĄŁŁłüŁŁõ ďŚŁŁðěŚŁŁĺŚŁŁœ ťc ŁŁÞŁŁÆŁŁì SS cŁŁþÃŁŁðýŚŁŁŻ ùÀŁŁł ÉŁŁ£ŁŁÈŁŁõ ćďŘÀŁŁœŚŁŁţŁŁĺŘ øć öŊ ďć ĆŁŁî ĆŁŁİöŊ

.ÀŁł ÀŁûŘĄŁą ÂŁÈŁţŁĂŁõ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ÂŁŞ üŁþŚŁûĆŁÞŁČŁÞŁő ,ćÀŁźŁõ ö cÀŁłďŚŁðěŚŁĺ ýŘÂŁŞ

ýŘÂŁŁŁŞ ÀŁŁŁîüŁŁŁœĄŁŁŁþ ěŘ ùćŚŁŁŁÔŁŁŁţŁŁŁĺŘ ŚŁŁŁŞ ýŘýďŚŁŁŁðěŚŁŁŁĺŚŁŁŁœ ĆŁŁŁœĄŁŁŁðžŁŁŁČŁŁŁû ćďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ SSŁŁŁþŘ ěŘ ýøďüŁŁŁŢ ŃŃŃ 2 ýďøŊćŚŁŁŁþ ŕĄŁÎŁą SSŁþŘ íc ÂŁţŁÈŁõ éc øÂŁ§ ,ùøĚŁäĆŁŞ .ÀŁĂŁîüŁÞŁœ ćŚŁźŁþŘ … ø ýćÂŁî ,øćďŘ ,üŁŞÂŁä ĆŁÜŁÞŁŻ ěŘ ÂŁÚŁþć ýŚŁûÍŁą üŁŁœŚŁŁÆŁŁØŁŁþ À cŁŁî ěŘ ŚŁŁûÍŁŁą SSŁŁþŘ ýĆŁŁÞŁŁû ďć ÓŁŁóŘ éc ÂŁŁ§ ,ñŚŁŁůŁŁõ ö cŘĄŁŁĂŁŁäĆŁŁŞ .ÀŁŁĂŁŁĂŁŁîüŁŁõ ùćŚŁŁÔŁŁţŁŁĺŘ íÂŁŁţŁŁÈŁŁõ ýŚŁŁûÀŁŁî ěŘ

.ÀŁĂŁîüŁõ ùćŚŁÔŁţŁĺŘ

üŁõŘÃŁóŘ â cŁŻŘÂŁõ 2

SSŁþÀŁŞ .ťŁĺŘ ùÀŁł ùćŘć áŚŁŻďŘ ŚŁûöŊ ĆŁŞ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ SS cŁţŁõ ďć ĆŁî ťŁĺŘ üŁŮŘďÂŁÖŁõ ý c øŚŁ§ ÂŁþě ü c ŁõŘÃŁóŘ íc ďŘÀŁõ ŚŁŁþ/ø ŤŚŁŁİ ¿ c ŁŁþďŚŁŁŮ ýŘďŘć â cŁŁŻŘÂŁŁõ ćcďĄŁŁõ ďć .ćĄŁŁłüŁŁõ ŠĄŁŁÆŁŁ¡ŁŁõ ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ ěŘ üŁŁŐÃŁŁŻ ,ŰŘďÂŁŁÖŁŁõ öŊ řŁŁČŁŁŮÂŁŁŮ ö cŘÂŁŞďŚŁî ťŁĺŘ ÂŁţŁúŁŞ ŘÁŁûâŁõ .ťŁÆŁČŁœ ÂŁÑŁœ ćcďĄŁõ íďŘÀŁõ SSŁþŘ ý c ÀŁãŁŞ ýŚŁûÂŁÑŁœÀ cŁþÀŁźŁŮ ø ŚŁûĆŁČŁ§ĚŁňŘ ,ÂŁÑŁœÀ cŁþÀŁźŁŮ ćcďĄŁŁõ Řď ÂŁŁþě ü c ŁŁõŘÃŁŁóŘ íc ďŘÀŁŁõ ýŚŁŁûÂŁŁÑŁŁœÀ cŁŁþÀŁŁźŁŁŮ ø ŚŁŁûĆŁŁČŁŁ§ĚŁŁňŘ SSŁŁþÂŁŁąŊ ćcÂŁŁŞďŚŁŁî ö cŚŁŁØŁŁõŘ ,ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ â cŁŁÔŁŁœýč öŊ Â cŁŁÑŁŁœÀ cŁŁþÀŁŁźŁŁŮ ŚŁŁþ/ø ŤŚŁŁİ SSŁŁþÂŁŁąŊ ,ÂŁŁÑŁŁœÀ cŁŁþÀŁŁźŁŁŮ ŚŁŁþ/ø ŤŚŁŁİ ¿ c ŁŁþďŚŁŁŮ ö cøÀŁŁŞ â cŁŁŻŘÂŁŁõ ćcďĄŁŁõ ďć .ÀŁŁĂŁŁûć ďŘÂŁŁì üŁŁĺďÂŁŁŞ

.ťŁĺŘ ÂŁÑŁœ ćcďĄŁõ ùÀŁł ùćŘć áŚŁŻďŘ ü c ŁõŘÃŁóŘ íc ďŘÀŁõ :ťŁĺŘ üŁõŘÃŁóŘ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ćcÂŁŞďŚŁî ýŘÂŁŞ ÂŁþě â cŁŻŘÂŁõ ěŘ ùćŚŁÔŁţŁĺŘ

1. The Unicode Consortium, The Unicode Standard, available from

http://www.unicode.org/

2. Mark Davis, “The Unicode Standard Annex #9, The Bidirectional Algorithm”,

available from http://www.unicode.org/unicode/reports/tr9/

3. Mark Davis, “The Unicode Standard Annex #13, Unicode Newline Guidelines”,

available from http://www.unicode.org/unicode/reports/tr13/

4. Mark Davis, Martin Dürst, “The Unicode Standard Annex #15, Unicode Normalization Forms”, available from

http://www.unicode.org/unicode/reports/tr15/

5. François Yergeau, “UTF-8, a transformation format of ISO 10646”, RFC 2279,

January 1998, available from http://www.ietf.org/rfc/rfc2279.txt

6. Paul Hoffman, François Yergeau, “UTF-16, an encoding of ISO 10646”, RFC 2781,

February 2000, available from http://www.ietf.org/rfc/rfc2781.txt

7. Mark Davis, “Unicode Standard Annex #19, UTF-32”,

available from http://www.unicode.org/unicode/reports/tr19/

ÓŁþďŚŁãŁŮ ø ŰŚŁ§ĚŁÎŁňŘ 3

:ćøďüŁõ ďŚŁîĆŁŞ ÂŁþě Ó c ŁþďŚŁãŁŮ ŚŁŞ ŚŁûùęŘø ŚŁþ/ø ŰŚŁ§ĚŁÎŁňŘ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ďć

SSŁţŁõ 1-3

text Â cŁŁŞŘÂŁŁŞ ďć .ťŁŁĺŘ ÂŁŁŮĄŁŁČŁŁşŁŁõŚŁŁî ýøď ùÀŁŁłùÂŁŁČŁŁąč ŚŁŁþ ùÀŁŁłÀŁŁî SS cŁŁţŁŁõ "SSŁŁţŁŁõ’ ěŘ ďĄŁŁÑŁŁĂŁŁõ fŚŁŁõĄŁŁÞŁŁä ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ ďć .ćøďüŁõ ďŚŁîĆŁŞ

ÍŁą 2-3

Â cŁŁŞŘÂŁŁŞ ďć .ćŘć öŚŁŁÈŁŁœ Řď öŚŁŁŞě ÀŁŁĂŁŁİ ŚŁŁþ ŒŁŁþ Ű cŚŁŁÞŁŁÜŁŁî öŊ ŚŁŁŞ öŘĄŁŁŮüŁŁõ ĆŁŁî ťŁŁĺŘ ŚŁŁûćŚŁŁÞŁŁœ ěŘ ýŘĆŁŁäĄŁŁÞŁŁźŁŁõ "ÍŁŁą’ .ćøďüŁõ ďŚŁîĆŁŞ script

ĆŁÆŁþĄŁœ 3-3

.ćøďüŁõ ďŚŁîĆŁŞ character Â cŁŞŘÂŁŞ ďć .öŊ Û cŁØŁł ěŘ ÛŁÖŁţŁÆŁõ ,ťŁĺŘ ùÀŁłĆŁţŁłĄŁœ SS cŁţŁõ À cŁ§Řø SSŁþÂŁţŁØŁİĄŁî "ĆŁÆŁþĄŁœ’

ĆŁÆŁþĄŁœĆŁäĄŁÞŁźŁõ 4-3

ďć .ÀŁœĄŁł ùćŚŁÔŁţŁĺŘ ýďŚŁţŁłĄŁœ Ű cŚŁäĚŁÏŘ ýĆŁŐŘďŘ ýŘÂŁŞ ĆŁî ťŁĺŘ üŁþŚŁûĆŁÆŁþĄŁœ ěŘ ýŘĆŁäĄŁÞŁźŁõ "ĆŁÆŁþĄŁœĆŁäĄŁÞŁźŁõ’

.ćøďüŁõ ďŚŁîĆŁŞ character set Â cŁŞŘÂŁŞ

ÛŁØŁł 5-3

ýćÀŁŁãŁŁţŁŁõ ýŚŁŁûÛŁŁØŁŁł ÀŁŁĂŁŁœŘĄŁŁŮüŁŁõ ŚŁŁûĆŁŁÆŁŁþĄŁŁœ .ťŁŁĺŘ ńŚŁŁą ýĆŁŁĂŁŁČŁŁõě ŒŁŁþ ďć ĆŁŁÆŁŁþĄŁŁœ üŁŁÈŁŁþŚŁŁÞŁŁœ Ű cďĄŁŁň "ÛŁŁØŁŁł’

.ćøďüŁõ ďŚŁîĆŁŞ glyph Â cŁŞŘÂŁŞ ďć .ÀŁĂŁłŚŁŞ ĆŁţŁłŘć

ùćŚŁĺ SS cŁţŁõ 6-3

ďŚŁŁŁîĆŁŁŁŞ plain text Â cŁŁŁŞŘÂŁŁŁŞ ďć .ťŁŁŁÆŁŁŁČŁŁŁœ ýŘĆŁŁŁŐŘďŘ ŚŁŁŁþ ýďŚŁŁŁţŁŁŁąŚŁŁŁĺ Ű cŚŁŁŁäĚŁŁŁÏŘ Û cŁŁŁõŚŁŁŁł ĆŁŁŁî ťŁŁŁĺŘ üŁŁŁĂŁŁŁţŁŁŁõ "ùćŚŁŁŁĺ SS cŁŁŁţŁŁŁõ’

.ćøďüŁõ

öćÂŁî ýďŘÁŁðÀŁî 7-3

.ćøďüŁõ ďŚŁîĆŁŞ encode Â cŁŞŘÂŁŞ ďć .ťŁĺŘ ŚŁûĆŁÆŁþĄŁœ ĆŁŞ ŚŁûÀŁî Œ c ŁþĆŁŞŒŁþ ń c ŚŁÊŁţŁąŘ "öćÂŁî ýďŘÁŁðÀŁî’

ŚŁûćŚŁÞŁœ 4

:ťŁĺŘ ùÀŁł ùćŚŁÔŁţŁĺŘ ÂŁþě ýŚŁûćŚŁÞŁœ ěŘ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ SS cŁţŁõ ďć

ďŚŁúŁİ ćcÀŁä ŒŁþ n öŊ ďć ĆŁî ,ćĄŁłüŁõ ÉŁ£ŁÈŁõ U+n Û cŁØŁł ĆŁŞ ÀŁîüŁœĄŁþ ýŚŁûĆŁÆŁþĄŁœ ŚŁŞ ÂŁÒŚŁĂŁţŁõ À cŁîe ŚŁþ ćÀŁä 10 SS cŁŁþÃŁŁðýŚŁŁŻ) F ŚŁŁŮ A SS cŁŁČŁŁŮĎ éc øÂŁŁ§ ø ,9 ŚŁŁŮ 0 SS cŁŁČŁŁŮĎ ô cŚŁŁìďŘ ěŘ ø ,ťŁŁĺŘ ùćÃŁŁœŚŁŁł ýŚŁŁĂŁŁśŁŁõ ďć üŁŁÞŁŁìďÇŁŁł ŚŁŁŮ .ÀŁŁłŚŁŁŞ ĆŁŁţŁŁłŘć ÝŁŁìď ďŚŁŁúŁŁİ ěŘ ÂŁŁţŁŁÞŁŁî ĆŁŁî SSŁŁþŘ ÂŁŁÚŁŁõ ,ćĄŁŁł áøÂŁŁł ÂŁŁÔŁŁň ŚŁŁŞ ÀŁŁþŚŁŁśŁŁœ n ćcÀŁŁä .ÀŁŁĂŁŁîüŁŁõ ùćŚŁŁÔŁŁţŁŁĺŘ (15 ŚŁŁŮ ťŁŁĺŘ SSŁŁØŁŁÞŁŁõ ŚŁŁûñøÀŁŁŻ ďć .U+102345 ø U+12345 ,U+1234 ,U+0123 ,U+0012 ,U+0001 :fĚŁŁůŁŁõ

.ćĄŁł éÁŁ§ U+ ďŚŁÊŁţŁąŘ ýŘÂŁŞ

.ťŁĺŘ "üŁĺďŚŁê Ã cŁČŁÞŁõ’ ýĆŁÆŁþĄŁœ À cŁîüŁœĄŁþ À cŁîeU+066B ŁŃŁŃ ñŚŁůŁõ

SSŁČŁóøŘ y ø x öŊ ďć ĆŁî ,ćĄŁłüŁõ ÉŁ£ŁÈŁõ x..y ŚŁþ U+x..U+y Û cŁØŁł ĆŁŞ ÀŁîüŁœĄŁþ ýŚŁûĆŁÆŁþĄŁœ ěŘ ýŘùćøÀŁ¡Łõ øć Û cŁŁŁõŚŁŁŁł ĆŁŁŁî ,ŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ěŘ ýŘĆŁŁŁţŁŁŁĺĄŁŁŁČŁŁŁŢ ýùćøÀŁŁŁ¡ŁŁŁõ Â cŁŁŁÚŁŁŁœŚŁŁŁþŚŁŁŁÞŁŁŁœ ŚŁŁŁûĆŁŁŁÎŁŁŁÖŁŁŁœ ø ÀŁŁŁœŘùćøÀŁŁŁ¡ŁŁŁõ ýŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ SSŁŁŁþÂŁŁŁąŊ ø

.ćĄŁłüŁõ ÃŁČŁœ ťŁĺÂŁúŁê Â cŁąŊ ø ñøŘ ýĆŁÆŁþĄŁœ

.ťŁĺŘ ÀŁîüŁœĄŁþ À cŁîe128 Û cŁõŚŁł U+0900..U+097F ŁŃŁŃ ñŚŁůŁõ

ÉŁ£ŁÈŁõ ⟨U+x, U+y, …, U+z⟩ \begin{matrix}Û \\ c \end{matrix}ŁØŁł ĆŁŞ ø ùÀŁł ŘÀŁŻ SSŁČŁŮĎ ñ cĄŁðÂŁþø ŚŁŞ ÀŁîüŁœĄŁþ À cŁîe ÀŁĂŁİ ŚŁþ øć ýĆŁóŚŁśŁœć

.ťŁĺŘ ýďŚŁČŁţŁąŘ U+ ćcŚŁÞŁœ .ťŁĺŘ ťŁĺŘď ĆŁŞ šŁİ ěŘ ëĄŁê ý c ďŘÁŁðćŚŁÞŁœ ďć ŚŁûĆŁÆŁþĄŁœ řc ŁČŁŮÂŁŮ .ćĄŁłüŁõ ŚŁŁûôŚŁŁœ SSŁŁþŘ .ťŁŁĺŘ ùćŘć ńŚŁŁÊŁŁţŁŁąŘ ŚŁŁţŁŁØŁŁþ üŁŁõŚŁŁœ ÀŁŁĂŁŁîüŁŁõ ÓŁŁþÂŁŁãŁŁŮ ĆŁŁî üŁŁþŚŁŁûĆŁŁÆŁŁþĄŁŁœ ýĆŁŁČŁŁÜŁŁî ĆŁŁŞ ÂŁŁőŚŁŁ§ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ ĆŁØŁÜŁŞ ,ÀŁĂŁţŁÆŁČŁœ ISO/IEC 10646 ø ÀŁîüŁœĄŁþ ýŚŁûćďŘÀŁœŚŁţŁĺŘ ýŚŁûĆŁÆŁþĄŁœ ü c ŁÆŁČŁÜŁÚŁœŘ ô cŚŁœ Õ cŁČŁìć ýĆŁÞŁŻÂŁŮ fŚŁõøÃŁó ø éøÂŁŁŁ§ ěŘ ÍŁŁŁÖŁŁŁê ŚŁŁŁûôŚŁŁŁœ SSŁŁŁþŘ ďć .ÀŁŁŁœŘùÀŁŁŁł ŠŚŁŁŁ£ŁŁŁţŁŁŁœŘ üŁŁŁĺďŚŁŁŁê ýŚŁŁŁûćÂŁŁŁŞďŚŁŁŁî ďć ŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ öŊ ćcÂŁŁŁŞďŚŁŁŁî ğ c ŚŁŁŁĺŘ ÂŁŁŁŞ

.ťŁĺŘ ùÀŁł ùćŚŁÔŁţŁĺŘ üŁĺďŚŁê ö cŚŁŞě ýŚŁûĆŁœŚŁÈŁœ

üŁĺďŚŁê ö cĄŁţŁõ ďć ùćŚŁÔŁţŁĺŘ ćcďĄŁõ ýŚŁûĆŁÆŁþĄŁœ 5

ùćŘć ÉŁŁČŁŁÊŁŁ£ŁŁŮ ŚŁŁûöŊ ĆŁŁŞ üŁŁÊŁŁ£ŁŁÈŁŁõ ýŚŁŁĂŁŁãŁŁõ ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ ďć ĆŁŁî ćÂŁŁČŁŁðüŁŁõ ÂŁŁŞ ďć Řď üŁŁþŚŁŁûĆŁŁÆŁŁþĄŁŁœ ÇŁŁ£ŁŁŞ SSŁŁþŘ ÂŁŁŞ fŚŁŁÖŁŁČŁŁìć Řď ĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ ÀŁŁþŚŁŁŞ ,ÀŁŁĂŁŁî üŁŁœŚŁŁśŁŁČŁŁţŁŁÈŁŁŢ ùÀŁŁõŊ ÇŁŁ£ŁŁŞ SSŁŁþŘ ďć ĆŁŁî ýŘĆŁŁÆŁŁþĄŁŁœ ěŘ ýćÂŁŁŞďŚŁŁî ÂŁŁðŘ .ťŁŁĺŘ ùÀŁŁł

.ÀŁĂŁî ÀŁČŁóĄŁŮ ŚŁþ ÂŁČŁÆŁÔŁŮ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ďć ùÀŁł ÂŁîč ýŚŁĂŁãŁõ ýŚŁĂŁśŁõ

ü c ŁœŚŁśŁČŁţŁÈŁŢ .ÀŁœŘùÀŁł ÉŁ£ŁÈŁõ ùďŚŁţŁĺ ťc ŁõĚŁä ŚŁŞ ĆŁî üŁþŚŁûĆŁÆŁþĄŁœ ÂŁÚŁõ ,ťŁĺŘ ýďŚŁśŁŻŘ ŚŁûĆŁÆŁþĄŁœ SSŁþŘ ü c ŁœŚŁśŁČŁţŁÈŁŢ ÂŁîč ýŚŁĂŁãŁõ ýŚŁĂŁśŁõ ÂŁŞ ÀŁþŚŁŞ ÃŁČŁœ ŚŁûöŊ ,öÀŁł üŁœŚŁśŁČŁţŁÈŁŢ Ű cďĄŁň ďć üŁóø ,ťŁĺŘ ýďŚŁČŁţŁąŘ ďŘćùďŚŁţŁĺ ýŚŁûĆŁÆŁþĄŁœ

.ÀŁœĄŁł ÀŁČŁóĄŁŮ ŚŁþ ÂŁČŁÆŁÔŁŮ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ďć ùÀŁł

ğ c ŚŁŁĺŘ ÂŁŁŞ fŚŁŁÖŁŁČŁŁìć ÀŁŁþŚŁŁŞ ùćŚŁŁÔŁŁţŁŁĺŘ SSŁŁþŘ ,ÀŁŁĂŁŁłŚŁŁŞ ĆŁŁţŁŁłŘć ÃŁŁČŁŁœ ýÂŁŁÚŁŁþć ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ĆŁŁŞ ěŚŁŁČŁŁœ ŚŁŁûćÂŁŁŞďŚŁŁî ĆŁŁî üŁŁŮďĄŁŁň ďć

.ćÂŁČŁð ŰďĄŁň ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ďć ùÀŁł ÓŁþÂŁãŁŮ ýŚŁĂŁãŁõ

ŚŁŁûĆŁŁÆŁŁþĄŁŁœ ĆŁŁŞ üŁŁÊŁŁ£ŁŁÈŁŁõ öŘÀŁŁĂŁŁİ ýŚŁŁĂŁŁãŁŁõ ISO/IEC 10646 ćcďŘÀŁŁœŚŁŁţŁŁĺŘ ĆŁŁî ŚŁŁŻöŊ ěŘ ŃŃŃ 1 ýďøŊćŚŁŁþ ŚŁŁŞ ÀŁŁþŚŁŁŞ ÂŁŁÚŁŁþć ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ěŘ ùćŚŁŁÔŁŁţŁŁĺŘ ø ťŁŁÆŁŁČŁŁœ üŁŁêŚŁŁî ćďŘÀŁŁœŚŁŁţŁŁĺŘ öŊ ŚŁŁŞ öćĄŁŁŞ ďŚŁŁðěŚŁŁĺ ,ÀŁŁûćüŁŁÞŁŁœ ÉŁŁČŁŁÊŁŁ£ŁŁŮ

.ÀŁłŚŁŞ ďŚŁðěŚŁĺ ÃŁČŁœ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ

ø ćďŘć üŁŮŚŁäĚŁÏŘ ýĆŁśŁĂŁŻ ÍŁÖŁê ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ďć ŚŁûĆŁÆŁþĄŁœ ýŘÂŁŞ ùÀŁłÉŁ£ŁÈŁõ Û cŁØŁł ŃŃŃ 2 ýďøŊćŚŁþ üŁÜŁØŁł ÂŁû ĆŁŞ ,ùćŚŁÔŁţŁĺŘ ćcďĄŁõ Ý cŁÜŁì ĆŁŞ ĆŁţŁÆŁŞ ÀŁœěŚŁźŁõ ŚŁûĆŁÆŁþĄŁœ ,ÀŁłŚŁŞ ùÀŁł ÂŁîč öŊ éc ĚŁą ĆŁî ýćďŘĄŁõ ďć ÂŁÚŁõ ěŘ ŚŁŁûĆŁŁÆŁŁþĄŁŁœ Ç c ŁŁþŚŁŁÞŁŁœ ýŘÂŁŁŞ ŚŁŁûćÂŁŁŞďŚŁŁî ťŁŁĺŘ SSŁŁØŁŁÞŁŁõ üŁŁţŁŁ§ .ÀŁŁœĄŁŁł ùćŘć ÇŁŁþŚŁŁÞŁŁœ ,ÀŁŁłŚŁŁŞ ĆŁŁÆŁŁþĄŁŁœ öŊ Â cŁŁÚŁŁœŚŁŁþŚŁŁÞŁŁœ ĆŁŁî

.ÀŁĂŁĂŁî ùćŚŁÔŁţŁĺŘ ćďŘÀŁœ üŁĺďŚŁê Íc Łą ĆŁŞ üŁţŁûŚŁśŁł ĆŁî ÛŁþÂ cŁŞ À cŁĂŁœŚŁõ üŁÎŁą

üŁóÂŁţŁĂŁî ýŚŁûĆŁÆŁþĄŁœ 1-5

üŁóÂŁţŁĂŁî ýŚŁûĆŁÆŁþĄŁœ ŃŃŃ 1 ñ cøÀŁŻ

ýďŚŁÊŁţŁąŘ ťc ŁõĚŁä  ĆŁÆŁþĄŁœ ô cŚŁœ  ÀŁîe
LF  ÀŁãŁŞ Â cŁÎŁĺ 000A
CR  ÂŁÎŁĺ Â cŁĺ 000D
ZWNJ  ýěŚŁźŁõ ýĆŁÜŁňŚŁê 200C
ZWJ  ýěŚŁźŁõ ñ cŚŁÊŁŮŘ 200D
LRM  ťŁĺŘďĆŁŞšŁİ ýĆŁœŚŁÈŁœ 200E
RLM  šŁİĆŁŞťŁĺŘď ýĆŁœŚŁÈŁœ 200F
LS  ŚŁûÂŁÎŁĺ ýùÀŁĂŁĂŁîŘÀŁŻ2028*
PS  ŚŁûÀŁĂŁŞ ýùÀŁĂŁĂŁîŘÀŁŻ 2029*
LRE  ťŁĺŘďĆŁŞšŁİ SS cŁţŁõÂŁþě 202A*
RLE  šŁİĆŁŞťŁĺŘď SS cŁţŁõÂŁþě 202B*
PDF  SSŁţŁõÂŁþě ö cŚŁþŚŁŢ 202C*
LRO  ťŁĺŘďĆŁŞšŁİ fŘÀŁČŁîŘ SS cŁţŁõÂŁþě 202D*
RLO  šŁİĆŁŞťŁĺŘď fŘÀŁČŁîŘ SS cŁţŁõÂŁþě 202E*
BOM  ŚŁûťŁþŚŁŞ řc ŁČŁŮÂŁŮ ýĆŁœŚŁÈŁœ FEFF

ÀŁŁþŚŁŁŞ ŚŁŁûÀŁŁĂŁŁŞ ø ŚŁŁûÂŁŁÎŁŁĺ ö cćÂŁŁîŘÀŁŁŻ ýŘÂŁŁŞ ,ÀŁŁłŚŁŁŞ UTF-8 řc ŁŁóŚŁŁì ďć SSŁŁţŁŁõ ĆŁŁî üŁŁŮďĄŁŁň ďć ŃŃŃ 1 ýďøŊćŚŁŁþ řc ŁŁŁóŚŁŁŁì ŚŁŁŁŞ ö cĄŁŁŁţŁŁŁõ ďć PS ø LS ěŘ ùćŚŁŁŁÔŁŁŁţŁŁŁĺŘ .ćĄŁŁŁł ùćŚŁŁŁÔŁŁŁţŁŁŁĺŘ ⟨CR, LF⟩ ŚŁŁŁþ ,CR ,LF ěŘ ćÂŁŁŁŞďŚŁŁŁî Â cŁŁŁţŁŁŁÆŁŁŁŞ ĆŁŁŁŞ ĆŁŁŁţŁŁŁÆŁŁŁŞ ĆŁŁãŁŁŻŘÂŁŁõ Ű ťc ŁŁĺĄŁŁČŁŁŢ ĆŁŁŞ ,ŚŁŁûĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ ěŘ ùćŚŁŁÔŁŁţŁŁĺŘ ĳ c ŁŁČŁŁ¡ŁŁň ýùĄŁŁČŁŁł ěŘ áĚŁŁÏŘ ýŘÂŁŁŞ .ťŁŁÆŁŁČŁŁœ ěŚŁŁźŁŁõ UTF-8

.ÀŁČŁĂŁî üŁŁĺďŚŁŁê ñ cŚŁŁÊŁŁŮŘ Ý cŁŁţŁŁþďĄŁŁÚŁŁóŘ ďć "ýěŚŁŁźŁŁõ ñ cŚŁŁÊŁŁŮŘ’ ø "ýěŚŁŁźŁŁõ ýĆŁŁÜŁŁňŚŁŁê’ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ŃŃŃ 2 ýďøŊćŚŁŁþ

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ Š ťc ŁĺĄŁČŁŢ ĆŁŞ ÂŁţŁÈŁČŁŞ Ű cŚŁäĚŁÏŘ ýŘÂŁŞ .ÀŁœøďüŁõ ďŚŁî ĆŁŞ

Ý cŁŁŁŁŁţŁŁŁŁŁþďĄŁŁŁŁŁÚŁŁŁŁŁóŘ ďć RLO ø ,LRO ,PDF ,RLE ,LRE ,RLM ,LRM ýŚŁŁŁŁŁûĆŁŁŁŁŁÆŁŁŁŁŁþĄŁŁŁŁŁœ ŃŃŃ 3 ýďøŊćŚŁŁŁŁŁþ

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ ÓóŘ ťc ŁĺĄŁČŁŢ ĆŁŞ ÂŁţŁÈŁČŁŞ Ű cŚŁäĚŁÏŘ ýŘÂŁŞ .ÀŁœøďüŁõ ďŚŁî ĆŁŞ ĆŁţŁúŁŻøć

ĆŁŁŁÆŁŁŁþĄŁŁŁœ SSŁŁŁþŘ ćĄŁŁŁŻø ô cÀŁŁŁä ŰďĄŁŁŁň ďć ĆŁŁŁî üŁŁŁœĄŁŁŁţŁŁŁõ ÃŁŁŁČŁŁŁČŁŁŁÞŁŁŁŮ ýŘÂŁŁŁŞ ÀŁŁŁþŚŁŁŁŞ BOM ýĆŁŁŁÆŁŁŁþĄŁŁŁœ ŃŃŃ 4 ýďøŊćŚŁŁŁþ ýŚŁŁŁŁûùÀŁŁŁŁœøÂŁŁŁŁŢ ýŘÀŁŁŁŁţŁŁŁŁŞŘ ďć ĆŁŁŁŁÆŁŁŁŁþĄŁŁŁŁœ SSŁŁŁŁþŘ ěŘ ùćŚŁŁŁŁÔŁŁŁŁţŁŁŁŁĺŘ .ćĄŁŁŁŁł ùćÂŁŁŁŁŞ ďŚŁŁŁŁî ĆŁŁŁŁŞ ,ÀŁŁŁŁœĄŁŁŁŁł ľěŘćÂŁŁŁŁŢ ùŚŁŁŁŁśŁŁŁŁţŁŁŁŁłŘ ťŁŁŁŁĺŘ SSŁŁŁŁØŁŁŁŁÞŁŁŁŁõ ćďŘÀŁœ üŁĂŁãŁõ ŚŁûťŁþŚŁŞ řŁČŁŮÂŁŮ ĆŁî UTF-8 ýŚŁûùÀŁœøÂŁŢ ýŘÀŁţŁŞŘ ďć üŁóø ćĄŁłüŁõ ĆŁČŁňĄŁŮ UTF-32 ø UTF-16 ĆŁŁŞ ÂŁŁţŁŁÈŁŁČŁŁŞ Ű cŚŁŁäĚŁŁÏŘ ýŘÂŁŁŞ .ťŁŁÆŁŁČŁŁœ ěŚŁŁźŁŁõ ÂŁŁÚŁŁþć À cŁŁňŚŁŁÖŁŁõ ýŘÂŁŁŞ ĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ ěŘ ùćŚŁŁÔŁŁţŁŁĺŘ .ćĄŁŁłüŁŁõ üŁŁúŁŁœ fŘÀŁŁþÀŁŁł

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ Ť ťc ŁĺĄŁČŁŢ

íÂŁţŁÈŁõ ý c ďŘÁŁðĆŁÎŁÖŁœ Ý cŁŐĚŁä 2-5

íÂŁţŁÈŁõ ý c ďŘÁŁðĆŁÎŁÖŁœ Ý cŁŐĚŁä ŃŃŃ 2 ñ cøÀŁŻ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ ÀŁîe
ĆŁÜŁňŚŁê 0020
.  ĆŁÎŁÖŁœ 002E
:  ĆŁÎŁÖŁœøć 003A
!  řŁźŁãŁŮ ťc ŁõĚŁä 0021
…  üŁÖŁêŘ ýĆŁÎŁÖŁœĆŁĺ 2026*
-  ùÂŁČŁŮ Íc Łą2010*
-  ŚŁúŁĂŁõùÂŁČŁŮ 002D
|  ýćĄŁÞŁä Íc Łą 007C
/  řŁþďŘ Íc Łą 002F
\  øďŘø řc ŁþďŘ Íc Łą 005C
∗  ùďŚŁţŁĺ 002A

íÂŁţŁÈŁõ ý c ďŘÁŁðĆŁÎŁÖŁœ Ý cŁŐĚŁä ŃŃŃ 2 ñ cøÀŁŻ ýĆŁõŘćŘ

üŁÈŁþŚŁÞŁœ Û cŁØŁł ĆŁÆŁþĄŁœ ô cŚŁœ ÀŁîe
)  ěŚŁŞ ÃŁţŁœŘÂŁŢ 0028
(  ĆŁţŁÆŁŞ ÃŁţŁœŘÂŁŢ 0029
[  ěŚŁŞ ĆŁłøÂŁî005B
]  ĆŁţŁÆŁŞ ĆŁłøÂŁî 005D
}  ěŚŁŞ ćĎĄŁî Ŋ 007B
{  ĆŁţŁÆŁŞ ćĎĄŁî Ŋ 007D
’  ěŚŁŞ ĆŁõĄŁČŁð 00AB
"  ĆŁţŁÆŁŞ ĆŁõĄŁČŁð00BB

ťc ŁŁõĚŁŁä ö cŘĄŁŁĂŁŁäĆŁŁŞ ŚŁŁþ ¿ŁŁþďŚŁŁŮ ýŘÃŁŁŻŘ ýěŚŁŁĺŘÀŁŁŻ ýŘÂŁŁŞ fŚŁŁŮÀŁŁÞŁŁä "řŁŁþďŘ Íc ŁŁą’ ýĆŁŁÆŁŁþĄŁŁœ ŃŃŃ 1 ýďøŊćŚŁŁþ :ćďŘć ŰøŚŁÔŁŮ (U+066B) "üŁĺďŚŁê Ã cŁČŁÞŁõ’ ŚŁŞ ĆŁÆŁþĄŁœ SSŁþŘ .ćøďüŁõ ďŚŁî ĆŁŞ SSŁţŁõ Û cŁąŘć ďć ÂŁÆŁî

1/4 = \frac{1}{4} = 0*25

1*4 = 1 + \frac{4}{10}

ĆŁŁî ýćďŘĄŁŁõ ďć ÂŁŁÚŁŁõ) ÀŁŁĂŁŁĂŁŁî ùćŚŁŁÔŁŁţŁŁĺŘ ĆŁŁÆŁŁþĄŁŁœ øć SSŁŁþŘ ýŘÂŁŁŞ üŁŁÔŁŁÜŁŁţŁŁ£ŁŁõ ü c ŁŁÈŁŁþŚŁŁÞŁŁœ ýŚŁŁûÛŁŁØŁŁł ěŘ ÀŁŁĂŁŁÔŁŁÒĄŁŁõ ŚŁŁûćÂŁŁŞďŚŁŁî

.(ùŘÂŁÞŁû SS cŁÔŁÜŁŮ ýŚŁûùŚŁÚŁţŁĺć À cŁĂŁœŚŁõ ,ćďŘć ćĄŁŻø üŁÈŁþŚŁÞŁœ ń c ŚŁą ýŚŁûťŁþćøÀŁ¡Łõ

,SSŁŁţŁŁõ ďć ćĄŁŁą ťc ŁŁČŁŁãŁŁìĄŁŁõ ĆŁŁŞ ĆŁŁţŁŁÆŁŁŞ ,ŚŁŁûŠĚŁŁì ø ŚŁŁûÃŁŁţŁŁœŘÂŁŁŢ Û cŁŁČŁŁśŁŁì ěŘ ,üŁŁţŁŁÔŁŁŻ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ŃŃŃ 2 ýďøŊćŚŁŁþ ďć ø ")’ Û cŁŁØŁŁłĆŁŁŞ šŁŁİĆŁŁŞťŁŁĺŘď ö cĄŁŁţŁŁõ ďć (U+0028) "ěŚŁŁŞ ÃŁŁţŁŁœŘÂŁŁŢ’ fĚŁŁůŁŁõ .ÀŁŁœÂŁŁþÁŁŁŢüŁŁõ üŁŁÔŁŁÜŁŁţŁŁ£ŁŁõ ýŚŁŁûÛŁŁØŁŁł

.ťŁĺŘ ùÀŁõŊ ÓóŘ ťc ŁĺĄŁČŁŢ ďć ďŚŁţŁêď SSŁþŘ đ c øÂŁÈŁõ .ćĄŁłüŁõ ÂŁûŚŁÒ "(’ Û cŁØŁłĆŁŞ ťŁĺŘďĆŁŞšŁİ öĄŁţŁõ

"ùÂŁŁŁČŁŁŁŮ Íc ŁŁŁą’ Œ c ŁŁŁČŁŁŁØŁŁŁÔŁŁŁŮ ĆŁŁŁî ćøďüŁŁŁõ ďŚŁŁŁî ĆŁŁŁŞ ýćďŘĄŁŁŁõ ďć ÍŁŁŁÖŁŁŁê "ŚŁŁŁúŁŁŁĂŁŁŁõùÂŁŁŁČŁŁŁŮ’ ýĆŁŁŁÆŁŁŁþĄŁŁŁœ ŃŃŃ 3 ýďøŊćŚŁŁŁþ ĆŁŁî ýÂŁŁÚŁŁþć řc ŁŁóŚŁŁì ěŘ ŚŁŁûùćŘć ĆŁŁî üŁŁõŚŁŁÚŁŁĂŁŁû fĚŁŁůŁŁõ ,ÀŁŁłŚŁŁśŁŁœ SSŁŁØŁŁÞŁŁõ (U+2212) "ŚŁŁúŁŁĂŁŁõ ťc ŁŁõĚŁŁä’ ěŘ (U+2010) ýŚŁĂŁãŁõ ö cćĄŁŞ ÉŁ£ŁÈŁõ Ű cďĄŁň ďć .ÀŁĂŁłŚŁŞ ùÀŁł ÛŁþÀŁśŁŮ ÀŁîüŁœĄŁþ řc ŁóŚŁì ĆŁŞ ÀŁœŘćüŁÞŁœ ÃŁþŚŁÞŁţŁõ Řď ĆŁÆŁþĄŁœ øć SSŁþŘ

.ćÂŁî ùćŚŁÔŁţŁĺŘ U+2212 ŚŁþ U+2010 üŁĂŁãŁþ ÕŁČŁìć ýŚŁûĆŁÆŁþĄŁœ ěŘ ÀŁþŚŁŞ ,ĆŁÆŁþĄŁœ

üŁĺďŚŁê ý c ďŘÁŁðĆŁÎŁÖŁœ Ý cŁŐĚŁä 3-5

üŁĺďŚŁê ý c ďŘÁŁðĆŁÎŁÖŁœ Ý cŁŐĚŁä ŃŃŃ 3 ñ cøÀŁŻ

üŁÈŁþŚŁÞŁœ Û cŁØŁł ĆŁÆŁþĄŁœ ô cŚŁœ ÀŁîe
,  üŁĺďŚŁê ñ cĄŁðÂŁþø060C
;  üŁĺďŚŁê ñ cĄŁðÂŁþøĆŁÎŁÖŁœ061B
?  üŁĺďŚŁê ñ cŘĄbŁĺ ťc ŁõĚŁä 061F
ŁŃ  üŁĺďŚŁê ü c ŁðÀŁČŁÈŁî 0640

üŁőŚŁþď Ý cŁŐĚŁä ø ôŚŁìďŘ 4-5

üŁőŚŁþď Ý cŁŐĚŁä ø ôŚŁìďŘ ŃŃŃ 4 ñ cøÀŁŻ

üŁÈŁþŚŁÞŁœ Û cŁØŁł ĆŁÆŁþĄŁœ ô cŚŁœ ÀŁîe
0  ÂŁÔŁň ü c ŁĺďŚŁê Ý cŁìď 06F0
1  ŒŁþ ü c ŁĺďŚŁê Ý cŁìď 06F1
2  øć ü c ŁĺďŚŁê Ý cŁìď 06F2
3  ĆŁĺ ü c ŁĺďŚŁê Ý cŁìď 06F3
4  ďŚŁúŁİ ü c ŁĺďŚŁê Ý cŁìď 06F4
5  ÿŁĂŁŢ ü c ŁĺďŚŁê Ý cŁìď06F5
ÇŁł ü c ŁĺďŚŁê Ý cŁìď 06F6
7  ťŁÔŁû ü c ŁĺďŚŁê Ý cŁìď 06F7
8  ťŁÈŁû ü c ŁĺďŚŁê Ý cŁìď 06F8
9  ĆŁœ ü c ŁĺďŚŁê Ý cŁìď 06F9
*  üŁĺďŚŁê Ã cŁČŁÞŁõ 066B
’  üŁĺďŚŁê ýŚŁûďŘÃŁû ýùÀŁĂŁĂŁîŘÀŁŻ 066C

üŁőŚŁþď Ý cŁŐĚŁä ø ôŚŁìďŘ ŃŃŃ 4 ñ cøÀŁŻ ýĆŁõŘćŘ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ ÀŁîe
%  üŁĺďŚŁê À cŁňďć 066A
+  ĆŁêŚŁőŘĆŁŞ ťc ŁõĚŁä002B
−  ŚŁúŁĂŁõ ťc ŁõĚŁä 2212*
×  ŠÂŁő ťc ŁõĚŁä 00D7
÷  ÝŁČŁÆŁÖŁŮ ťc ŁõĚŁä00F7*
<  ÂŁţŁØŁİĄŁî ťc ŁõĚŁä 003C
=  ýøŚŁÆŁõ ťc ŁõĚŁä 003D
>  ÂŁţŁðďÃŁŞ ťc ŁõĚŁä003E

ýŚŁûÛŁØŁł ,SSŁţŁõ ďć ćĄŁą ťc ŁČŁãŁìĄŁõ ĆŁŞ ĆŁţŁÆŁŞ ,"ÂŁţŁðďÃŁŞ ťc ŁõĚŁä’ ø "ÂŁţŁØŁİĄŁî ťc ŁõĚŁä’ ŃŃŃ ýďøŊćŚŁþ ŚŁŞ ëĄŁê ñ cøÀŁŻ ďć ŚŁûĆŁÆŁþĄŁœ SSŁþŘ Û cŁØŁł .ťŁĺŘ ùÀŁõŊ ÓóŘ ťc ŁĺĄŁČŁŢ ďć ďŚŁţŁêď SSŁþŘ đ c øÂŁÈŁõ .ÀŁœÂŁČŁðüŁõ üŁÔŁÜŁţŁ£Łõ

.ťŁĺŘ ùÀŁõŊ ,ćŘÀŁäŘ ö cŚŁČŁõ ďć üŁĂŁãŁþ ,ŚŁûöŊ ñ cĄŁÞŁãŁõ ýĆŁĂŁČŁõě ĆŁŞ ĆŁŻĄŁŮ

üŁĺďŚŁê ü c ŁÜŁňŘ éc øÂŁ§ 5-5

üŁĺďŚŁê ü c ŁÜŁňŘ éc øÂŁ§ ŃŃŃ 5 ñ cøÀŁŻ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ  ÀŁîe
Ŕ  ùÃŁÞŁû ü c ŁĺďŚŁê éc ÂŁ§0621
Ŋ  Ŋ ü c ŁĺďŚŁê éc ÂŁ§ 0622
Ř  ÓŁóŘ ü c ŁĺďŚŁê éc ÂŁ§0627
Řb  ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ ü c ŁĺďŚŁê éc ÂŁ§ 0623
Š  Š c ü c ŁĺďŚŁê éc ÂŁ§ 0628
Ť  Ť c ü c ŁĺďŚŁê éc ÂŁ§ 067E

üŁĺďŚŁê ü c ŁÜŁňŘ éc øÂŁ§ ŃŃŃ 5 ñ cøÀŁŻ ýĆŁõŘćŘ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ ÀŁîe
Ű  Ű c ü c ŁĺďŚŁê éc ÂŁ§ 062A
Ÿ  Ÿ c ü c ŁĺďŚŁê éc ÂŁ§ 062B
Ž  ÝŁČŁŻ ü c ŁĺďŚŁê éc ÂŁ§062C
Ĳ  Ĳ c ü c ŁĺďŚŁê éc ÂŁ§ 0686
đ  đ c ü c ŁĺďŚŁê éc ÂŁ§ 062D
ă  ă c ü c ŁĺďŚŁê éc ÂŁ§ 062E
ć  ñŘć ü c ŁĺďŚŁê éc ÂŁ§ 062F
č  ñŘč ü c ŁĺďŚŁê éc ÂŁ§ 0630
ď  ď cü c ŁĺďŚŁê éc ÂŁ§ 0631
ě  ěc ü c ŁĺďŚŁê éc ÂŁ§ 0632
ę  ęc ü c ŁĺďŚŁê éc ÂŁ§ 0698
ğ  SSŁČŁĺ ü c ŁĺďŚŁê éc ÂŁ§ 0633
ľ  SSŁČŁł ü c ŁĺďŚŁê éc ÂŁ§0634
ń  ćŚŁň ü c ŁĺďŚŁê éc ÂŁ§ 0635
ŋ  ćŚŁő ü c ŁĺďŚŁê éc ÂŁ§ 0636
ŕ  ŚŁÏ ü c ŁĺďŚŁê éc ÂŁ§ 0637
à  ŚŁÒ ü c ŁĺďŚŁê éc ÂŁ§ 0638
á  SSŁČŁä ü c ŁĺďŚŁê éc ÂŁ§0639
å  SSŁČŁè ü c ŁĺďŚŁê éc ÂŁ§063A
é  éc ü c ŁĺďŚŁê éc ÂŁ§ 0641
ë  éŚŁì ü c ŁĺďŚŁê éc ÂŁ§ 0642

üŁĺďŚŁê ü c ŁÜŁňŘ éc øÂŁ§ ŃŃŃ 5 ñ cøÀŁŻ ýĆŁõŘćŘ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ  ÀŁîe
í  éŚŁî ü c ŁĺďŚŁê éc ÂŁ§ 06A9
ï  éŚŁð ü c ŁĺďŚŁê éc ÂŁ§ 06AF
ñ  ôĎ ü c ŁĺďŚŁê éc ÂŁ§ 0644
ô  ÝŁČŁõ ü c ŁĺďŚŁê éc ÂŁ§0645
ö  öĄŁœ ü c ŁĺďŚŁê éc ÂŁ§ 0646
ø  øŘø ü c ŁĺďŚŁê éc ÂŁ§0648
øb  ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ øŘø ü c ŁĺďŚŁê éc ÂŁ§ 0624
Łû  Łû c ü c ŁĺďŚŁê éc ÂŁ§ 0647
ý ý c ü c ŁĺďŚŁê éc ÂŁ§ 06CC
ý b  ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ý c ü c ŁĺďŚŁê éc ÂŁ§ 0626

.ćÂŁŁî ĆŁŁóćŚŁŁśŁŁõ ÃŁŁČŁŁœ ĆŁŁÆŁŁþĄŁŁœ øć Ű cďĄŁŁňĆŁŁŞ öŘĄŁŁŮüŁŁõ Řď ëĄŁŁê ñøÀŁŁŻ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ěŘ üŁŁÌŁŁãŁŁŞ ŃŃŃ 1 ýďøŊćŚŁŁþ ĆŁóćŚŁśŁõ ⟨U+0627, U+0653⟩ \begin{matrix}Ű \\ c \end{matrix}ďĄŁňĆŁŞ ÝŁû ø U+0622 Ű cďĄŁňĆŁŞ ÝŁû öŘĄŁŮüŁõ Řď "Ŋ ü c ŁĺďŚŁê éc ÂŁ§’ fĚŁůŁõ Ÿ ťc ŁŁŁĺĄŁŁŁČŁŁŁŢ ĆŁŁŁŞ ,ùďŚŁŁŁŞ SSŁŁŁþŘ ďć ÂŁŁŁţŁŁŁÖŁŁŁČŁŁŁìć á c ĚŁŁŁÏŘ ýŘÂŁŁŁŞ .ťŁŁŁĺŘ ĳŁŁŁŻÂŁŁŁõ ýŘĆŁŁŁÆŁŁŁþĄŁŁŁœŒŁŁŁŮ Û cŁŁŁØŁŁŁł ,ćďŘĄŁŁŁõ SSŁŁŁþŘ ďć .ćÂŁŁŁî

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ

"SSŁČŁä ü c ŁĺďŚŁê éc ÂŁ§’ fĚŁůŁõ ,ÀŁœÂŁþÁŁŢüŁõ ćĄŁą ĆŁŞ üŁÔŁÜŁţŁ£Łõ ýŚŁûÛŁØŁł üŁĺďŚŁê éc øÂŁ§ ŃŃŃ 2 ýďøŊćŚŁþ SSŁŁŁČŁŁŁČŁŁŁãŁŁŁŮ üŁŁŁĺďŚŁŁŁê ñ cŚŁŁŁÊŁŁŁŮŘ Ý cŁŁŁţŁŁŁþďĄŁŁŁÚŁŁŁóŘ ďć ŚŁŁŁûÛŁŁŁØŁŁŁł SSŁŁŁþŘ .ćĄŁŁŁłüŁŁŁõ ùÀŁŁŁþć "á’ ø "â’ ,"ŁŁŁã’ ,"ŁŁŁä’ ýŚŁŁŁûÛŁŁŁØŁŁŁł ĆŁŁŁŞ ÍŁŁÖŁŁê ÀŁŁœŘùÀŁŁõŊ ëĄŁŁê ñ cøÀŁŁŻ ďć ĆŁŁî üŁŁþŚŁŁûÛŁŁØŁŁł .ťŁŁĺŘ ùÀŁŁł ĳŁŁþÂŁŁÈŁŁŮ Š ťc ŁŁĺĄŁŁČŁŁŢ ďć ÝŁŁţŁŁþďĄŁŁÚŁŁóŘ SSŁŁþŘ .ÀŁŁœĄŁŁłüŁŁõ

.ÀŁœďŘć üŁŮŚŁäĚŁÏŘ ýĆŁśŁĂŁŻ

üŁäÂŁê éc øÂŁ§ 6-5

üŁäÂŁê éc øÂŁ§ ŃŃŃ 6 ñ cøÀŁŻ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ  ÀŁîe

SSŁČŁþŚŁŢ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ éc ÂŁ§

                                                0625*

ÛŁňø Ó c ŁóŘ éc ÂŁ§ 0671*
üŁŞÂŁä éc ŚŁî éc ÂŁ§ 0643*
Ę  ćÂŁð Ű c éc ÂŁ§ 0629

ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c éc ÂŁ§

                                                064A*

ý  ĆŁÎŁÖŁœüŁŞ ü c ŁŞÂŁä ý c éc ÂŁ§0649*

éc ÂŁ§’ ěŘ ùćŚŁÔŁţŁĺŘ ø "éŚŁî ü c ŁĺďŚŁê éc ÂŁ§’ ýŚŁŻ ĆŁŞ "üŁŞÂŁä éc ŚŁî éc ÂŁ§’ ěŘ ùćŚŁÔŁţŁĺŘ ŃŃŃ 1 ýďøŊćŚŁþ .ťŁÆŁČŁœ ěŚŁźŁõ öŘĄŁĂŁä žŁČŁûĆŁŞ "ý ü c ŁĺďŚŁê éc ÂŁ§’ ýŚŁŻ ĆŁŞ "ĆŁÎŁÖŁœüŁŞ ü c ŁŞÂŁä ý c éc ÂŁ§’ ŚŁþ "ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c ĆŁŁŞ SSŁŁţŁŁõ ŚŁŁþ ,ùćĄŁŁŞ ÂŁŁÑŁŁœ ćďĄŁŁõ ŚŁŁûöŊ ń c ŚŁŁą Û cŁŁØŁŁł ĆŁŁî ćÂŁŁî ùćŚŁŁÔŁŁţŁŁĺŘ ŚŁŁûĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ ěŘ öŘĄŁŁŮüŁŁõ üŁŁŮďĄŁŁň ďć ŚŁŁúŁŁĂŁŁŮ ÇŁŁþŚŁŁÞŁŁœ ŚŁŁûöŊ ťc ŁŁĺďć Û cŁŁØŁŁł ĆŁŁŞ Řď ŚŁŁûĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ ÀŁŁĂŁŁÔŁŁÒĄŁŁõ ŚŁŁûćÂŁŁŞďŚŁŁî .ÀŁŁłŚŁŁŞ ŚŁŁûöŊ ñ cŚŁŁůŁŁõŘ ø øćďŘ ,üŁŁŞÂŁŁä ö cŚŁŁŞě

.ćĄŁł ùćŘć ÇŁþŚŁÞŁœ ĆŁÎŁÖŁœ ö cøÀŁŞ ÀŁþŚŁśŁœ ùŚŁðžŁČŁû "ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c éc ÂŁ§’ .ÀŁĂŁûć

ø ÂŁąŊ ýŚŁûÛŁØŁł ďć ŚŁõŘ ,ťŁĺŘ "éŚŁî ü c ŁĺďŚŁê éc ÂŁ§’ ÀŁĂŁœŚŁõ ÍŁĺø ø ñøŘ ýŚŁûÛŁØŁł ďć "üŁŞÂŁä éc ŚŁî éc ÂŁ§’ ýŚŁûÛŁØŁł ďć "ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c éc ÂŁ§’ .ćĄŁłüŁõ ÂŁûŚŁÒ ùÃŁÞŁû Ć cŁČŁśŁł üŁţŁõĚŁä ù cŘÂŁÞŁû ĆŁŞ ø ÇŁîÂŁĺ ö cøÀŁŞ ŚŁúŁĂŁŮ ÂŁŁŁûŚŁŁŁÒ ľÂŁŁŁþě ďć ĆŁŁŁÎŁŁŁÖŁŁŁœøć ŚŁŁŁŞ ŚŁŁŁúŁŁŁĂŁŁŁŮ ø ÂŁŁŁąŊ ýŚŁŁŁûÛŁŁŁØŁŁŁł ďć ŚŁŁŁõŘ ,ťŁŁŁĺŘ "ý c ü c ŁŁŁĺďŚŁŁŁê éc ÂŁŁŁ§’ ÀŁŁŁĂŁŁŁœŚŁŁŁõ ÍŁŁŁĺø ø ñøŘ ďć ŚŁŁõŘ ,ťŁŁĺŘ "ý c ü c ŁŁĺďŚŁŁê éc ÂŁŁ§’ ÀŁŁĂŁŁœŚŁŁõ ŚŁŁúŁŁĂŁŁŮ ø ÂŁŁąŊ ýŚŁŁûÛŁŁØŁŁł ďć "ĆŁŁÎŁŁÖŁŁœüŁŁŞ ü c ŁŁŞÂŁŁä ý c éc ÂŁŁ§’ .ćĄŁŁłüŁŁõ

.ćĄŁłüŁõ ÂŁûŚŁÒ ĆŁÎŁÖŁœ ö cøÀŁŞ ÍŁĺø ø ñøŘ ýŚŁûÛŁØŁł

üŁĺďŚŁê ýŚŁûĆŁœŚŁÈŁœ 7-5

üŁĺďŚŁê ýŚŁûĆŁœŚŁÈŁœ ŃŃŃ 7 ñ cøÀŁŻ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ  ÀŁîe
(ĆŁ¡ŁţŁê) üŁĺďŚŁê Â cŁŞdě 064E
(ùÂŁÆŁî) üŁĺďŚŁê Â cŁþě 0650
(ĆŁÞŁő) üŁĺďŚŁê Ç c ŁČŁŢ 064F
(řŁÊŁœ SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Â cŁŞděøć064B

(ÂŁŻ SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Â cŁþěøć

                                                 064D

(âŁêď SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Ç c ŁČŁŢøć

                                                 064C

üŁĺďŚŁê À cŁþÀŁÈŁŮ 0651
üŁĺďŚŁê SS cŁîŚŁĺ 0652
üŁĺďŚŁê À cŁõ 0653*
ĎŚŁŞ ü c ŁĺďŚŁê ýùÃŁÞŁû0654

SSŁČŁþŚŁŢ ü c ŁĺďŚŁê ýùÃŁÞŁû

                                                 0655*

üŁĺďŚŁê ýùďĄŁÊŁÖŁõ Ó c ŁóŘ0670

Â cŁŁþě ŚŁŁþ ĎŚŁŁŞ ,ćďĄŁŁõ řc ŁŁÆŁŁ§ ÂŁŁŞ ÀŁŁþŚŁŁŞ ø ÀŁŁœďŘć ùÀŁŁœĄŁŁłřŁŁČŁŁîÂŁŁŮ ťc ŁŁČŁŁňŚŁŁą ëĄŁŁê ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ŃŃŃ 1 ýďøŊćŚŁŁþ Š ťc ŁŁĺĄŁŁČŁŁŢ ĆŁŁŞ ,ñŚŁŁÊŁŁŮŘ Ý cŁŁţŁŁþďĄŁŁÚŁŁóŘ ÂŁŁŞ ŚŁŁûĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ Â cŁŁČŁŁŹŚbŁŁŮ ćcďĄŁŁõ ďć .ÀŁŁœĄŁŁł ùćŘć ÇŁŁþŚŁŁÞŁŁœ ćĄŁŁą ěŘ ÛŁŁśŁŁì ýĆŁŁÆŁŁþĄŁŁœ

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ

ýøď "SSŁČŁþŚŁŢ ü c ŁĺďŚŁê ýùÃŁÞŁû’ ø "ĎŚŁŞ ü c ŁĺďŚŁê ýùÃŁÞŁû’ ýŚŁûĆŁÆŁþĄŁœ ĆŁî üŁŮďĄŁň ďć ŃŃŃ 2 ýďøŊćŚŁþ ťŁŁĺć ěŘ Řď ćĄŁŁą ýŚŁŁûĆŁŁÎŁŁÖŁŁœ üŁŁĺÂŁŁî ýĆŁŁÆŁŁþĄŁŁœ ,ÀŁŁĂŁŁþŚŁŁČŁŁŞ "ďŘćĆŁŁÎŁŁÖŁŁœ ü c ŁŁŞÂŁŁä ý c éc ÂŁŁ§’ ŚŁŁþ ,"ý c ü c ŁŁĺďŚŁŁê éc ÂŁŁ§’

.ÀŁûćüŁõ

ùćŚŁŁÔŁŁţŁŁĺŘ ńŚŁŁą ýŚŁŁûÛŁŁØŁŁł ěŘ ŚŁŁûĆŁŁœŚŁŁÈŁŁœ řc ŁŁČŁŁîÂŁŁŮ Ç c ŁŁþŚŁŁÞŁŁœ ýŘÂŁŁŞ ÀŁŁĂŁŁœŘĄŁŁŮüŁŁõ ŚŁŁûćÂŁŁŞďŚŁŁî ŃŃŃ 3 ýďøŊćŚŁŁþ "üŁĺďŚŁê Â cŁþě’ ö cćŘć ÇŁþŚŁÞŁœ ýŚŁŻ ĆŁŞ ťŁĺŘ ÂŁţŁúŁŞ "üŁĺďŚŁê Â cŁþě’ ø "üŁĺďŚŁê À cŁþÀŁÈŁŮ’ řc ŁČŁîÂŁŮ ýŘÂŁŞ fĚŁůŁõ .ÀŁĂŁĂŁî

.ćŘć ÇŁþŚŁÞŁœ "üŁĺďŚŁê À cŁþÀŁÈŁŮ’ Â cŁþě ďć Řď öŊ ,üŁĺÂŁî éc ÂŁ§ Â cŁþě ďć

áĄŁĂŁÞŁõ ýŚŁûĆŁÆŁþĄŁœ 8-5

ø üŁŞÂŁä À cŁĂŁœŚŁõ ÂŁÚŁþć ýŚŁûöŚŁŞě ö cĄŁţŁõ ďć ŚŁûöŊ ěŘ ùćŚŁÔŁţŁĺŘ .ÀŁœĄŁł ùćŚŁÔŁţŁĺŘ üŁĺďŚŁê ö cĄŁţŁõ ďć ÀŁþŚŁśŁœ ŚŁûĆŁÆŁþĄŁœ SSŁþŘ .ćÂŁČŁð ŰďĄŁň ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ďć ćĄŁŻĄŁõ Ó c ŁþÂŁãŁŮ ğ c ŚŁĺŘ ÂŁŞ ÀŁþŚŁŞ øćďŘ

áĄŁĂŁÞŁõ ýŚŁûĆŁÆŁþĄŁœ ŃŃŃ 8 ñ cøÀŁŻ

üŁÈŁþŚŁÞŁœ Û cŁØŁł  ĆŁÆŁþĄŁœ ô cŚŁœ  ÀŁîe
ùb  ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ øćďŘ ù céc ÂŁ§ 06C0
0  üŁŞÂŁä Â cŁÔŁň Ý cŁìď 0660
1  üŁŞÂŁä Œ c Łþ Ý cŁìď 0661
2  üŁŞÂŁä ø cć Ý cŁìď 0662
3  üŁŞÂŁä Ć cŁĺ Ý cŁìď 0663
üŁŞÂŁä ď cŚŁúŁİ Ý cŁìď 0664

üŁŞÂŁä ÿ c ŁĂŁŢ Ý cŁìď

                                                 0665

6  üŁŞÂŁä Ç c Łł Ý cŁìď 0666
7  üŁŞÂŁä ťc ŁÔŁû Ý cŁìď 0667
8  üŁŞÂŁä ťc ŁÈŁû Ý cŁìď 0668
9  üŁŞÂŁä Ć cŁœ Ý cŁìď 0669

SSŁŁþŘ .ćďŘć üŁŁŮŚŁŁäĚŁŁÏŘ ýĆŁŁśŁŁĂŁŁŻ Íc ŁŁÖŁŁê ø ťŁŁÆŁŁČŁŁœ ćďŘÀŁŁœŚŁŁţŁŁĺŘ ëĄŁŁê ñ cøÀŁŁŻ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ô cŚŁŁœ ŃŃŃ 1 ýďøŊćŚŁŁþ

.ÀŁûćüŁÞŁœ ńŚŁÊŁţŁąŘ üŁõŚŁœ ëĄŁê ñ cøÀŁŻ ýŚŁûĆŁÆŁþĄŁœ ĆŁŞ ćďŘÀŁœŚŁţŁĺŘ

ýŘÂŁŁŞ .ćĄŁŁł ùćŚŁŁÔŁŁţŁŁĺŘ üŁŁĺďŚŁŁê ö cĄŁŁţŁŁõ ýŘÂŁŁŞ öŘĄŁŁĂŁŁä žŁŁČŁŁû ĆŁŁŞ ÀŁŁþŚŁŁśŁŁœ U+06C0 ýĆŁŁÆŁŁþĄŁŁœ ŃŃŃ 2 ýďøŊćŚŁŁþ ýùÃŁŁŁÞŁŁŁû’ ù cŘÂŁŁŁÞŁŁŁûĆŁŁŁŞ (U+0647) "ŁŁŁû c üŁŁŁĺďŚŁŁŁê éc ÂŁŁŁ§’ ýĆŁŁŁÆŁŁŁþĄŁŁŁœ ěŘ ÀŁŁŁþŚŁŁŁŞ "ŚŁŁŁõ ĆbŁŁŁœŚŁŁŁą’ Û cŁŁŁůŁŁŁõ üŁŁŁŮŘďŚŁŁŁśŁŁŁä SS cŁŁŁţŁŁŁłĄŁŁŁœ ĆŁÆŁþĄŁœ SSŁþŘ ĆŁŞ üŁĺďŚŁê ö cĄŁţŁõ ďć ĆŁî üŁŮďĄŁň ďć ÀŁĂŁœŘĄŁŮüŁõ ŚŁûćÂŁŞďŚŁî .ćĄŁł ùćŚŁÔŁţŁĺŘ (U+0654) "ĎŚŁŞ ü c ŁĺďŚŁê ĆŁî üŁŮďĄŁň ďć .ÀŁĂŁĂŁî SSŁþÃŁðýŚŁŻ ⟨0647, 0654, 200C⟩ ŚŁþ ⟨0647, 0654⟩ ŚŁŞ ćďĄŁõ ĆŁŞ ĆŁţŁÆŁŞ Řď öŊ ÀŁœćďĄŁąÂŁŞ

.ćÂŁČŁÚŁŞ ŰďĄŁň ÀŁþŚŁśŁœ üŁĂŁþÃŁðýŚŁŻ SSŁþŘ ,ÀŁłŚŁśŁœ üŁĺďŚŁê ö cŚŁŞě ĆŁŞ SSŁţŁõÂŁþě ŚŁþ SSŁţŁõ ÀŁûŘĄŁ£ŁŞ ćÂŁŞďŚŁî ĆŁî ýćďŘĄŁõ ďć ÂŁÚŁõ (U+0660..U+0669) üŁŞÂŁä ô cŚŁìďŘ ěŘ ùćŚŁÔŁţŁĺŘ ŃŃŃ 3 ýďøŊćŚŁþ üŁœŚŁśŁČŁţŁÈŁŢ Řď üŁŞÂŁä ô cŚŁìďŘ ÀŁĂŁûŘĄŁ£ŁŞ ĆŁî üŁþŚŁûćÂŁŞďŚŁî .ťŁÆŁČŁœ ěŚŁźŁõ ćĄŁł ÛŁŐŚŁì ÃŁþŚŁÞŁŮ üŁŞÂŁä ø üŁĺďŚŁê ô cŚŁìďŘ ö cŚŁČŁõ ĆŁî ćĄŁł ťŁìć ÀŁþŚŁŞ .ÀŁœĄŁł ÛŁŐŚŁì ÃŁþŚŁÞŁŮ üŁŞÂŁä ø üŁĺďŚŁê Ç c Łł ø ,ÿŁĂŁŢ ,ďŚŁúŁİ ô cŚŁìďŘ Û cŁØŁł ö cŚŁČŁõ ťŁĺŘ ÂŁţŁúŁŞ ÀŁĂŁĂŁî

.ÀŁœďŘć ŰøŚŁÔŁŮ ÀŁîüŁœĄŁþ ýĆŁţŁúŁŻøć Ý cŁţŁþďĄŁÚŁóŘ ďć ýÂŁþÁŁŢťŁúŁŻ à c ŚŁ¡Łó ěŘ üŁŞÂŁä ø üŁĺďŚŁê ô cŚŁìďŘ

ăĄŁÆŁĂŁõ ýŚŁûĆŁÆŁþĄŁœ 9-5

ÃŁŁČŁŁœ ćďŘÀŁŁœŚŁŁţŁŁĺŘ SSŁŁþŘ ďć ,ÀŁŁœŘùÀŁŁł ÉŁŁ£ŁŁÈŁŁõ ăĄŁŁÆŁŁĂŁŁõ ö cŘĄŁŁĂŁŁäĆŁŁŞ ÀŁŁîüŁŁœĄŁŁþ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ ďć ĆŁŁî üŁŁþŚŁŁûĆŁŁÆŁŁþĄŁŁœ ýĆŁŁČŁŁÜŁŁî ŚŁŁŁûöŊ ĆŁŁŁŞ öćďĄŁŁŁąÂŁŁŁŞ Ű cďĄŁŁŁň ďć ø ,ÀŁŁŁĂŁŁŁĂŁŁŁî ÀŁŁŁČŁŁŁóĄŁŁŁŮ Řď ŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ SSŁŁŁþŘ ÀŁŁŁþŚŁŁŁśŁŁŁœ ŚŁŁŁûćÂŁŁŁŞďŚŁŁŁî .ÀŁŁŁœĄŁŁŁłüŁŁŁõ üŁŁŁÖŁŁŁÜŁŁŁŮ ăĄŁŁŁÆŁŁŁĂŁŁŁõ

.ÀŁĂŁĂŁî ÂŁČŁÆŁÔŁŮ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ýĆŁ£ŁÆŁœ SSŁþÂŁąŊ ğ c ŚŁĺŘÂŁŞ Řď ŚŁûöŊ ŚŁþ ,ùćÂŁî üŁłĄŁŢÝŁÈŁİ ŚŁûöŊ ěŘ ÀŁĂŁœŘĄŁŮüŁõ

ÓóŘ ťc ĺĄČŢ

ĆŁţŁúŁŻøć Ý cŁţŁþďĄŁÚŁóŘ

(üŁõŘÃŁóŘ)

Ý cŁŐĚŁä ø ćŘÀŁäŘ Û cŁõŚŁł ü c ŁĺďŚŁê ö cĄŁţŁõ ĆŁî ŚŁŻöŊ ěŘ ø ,SSŁČŁŮĎ ø üŁĺďŚŁê ýŚŁûÍŁą SS cŁţŁłĄŁœ ťc ŁúŁŻ Ű cøŚŁÔŁŮ ťc ŁÜŁä ĆŁŞ SSŁþŘ ýŚŁûĆŁÆŁþĄŁœ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ,ÀŁœĄŁłüŁõ øďĆŁŞøď ôŚŁúŁŞŘ ŚŁŞ ľěŘćÂŁŢ ô cŚŁÚŁĂŁû ďć ,ĆŁœŚŁŞěÀŁĂŁİ ö cĄŁţŁõ ŚŁþ ,üŁőŚŁþď .ÀŁĂŁîüŁõ ýďŘÁŁðÀŁî ÀŁœďÁŁðüŁõ SSŁţŁõ ýùÀŁĂŁœŘĄŁą SS cŁûč ěŘ ŚŁûĆŁÆŁþĄŁœ ĆŁî üŁśŁČŁŮÂŁŮ üŁĂŁãŁþ ,üŁþŚŁĂŁãŁõ řc ŁČŁŮÂŁŮ ĆŁŞ Řď öĄŁţŁõ

.ćøďüŁõ ďŚŁîĆŁŞ ÇŁþŚŁÞŁœ Û cŁŞŚŁì řc ŁČŁŮÂŁŮ ŒŁþ ĆŁŞ řŁČŁŮÂŁŮ SSŁþŘ Û cŁþÀŁśŁŮ ýŘÂŁŞ ĆŁţŁúŁŻøć Ý cŁţŁþďĄŁÚŁóŘ

Ç c ŁþŚŁÞŁœ ý c ŘÂŁŞ SSŁþŘÂŁŞŚŁĂŁŞ .ÀŁœĄŁł ĆŁóćŚŁśŁõ üŁþŚŁĂŁãŁõ řc ŁČŁŮÂŁŮ ĆŁŞ ÀŁþŚŁŞ ŚŁûĆŁÆŁþĄŁœ ,ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ÂŁŞ üŁĂŁţŁśŁõ ýŚŁûćÂŁŞďŚŁî ďć .ćĄŁŁŁł ÛŁŁŁþÀŁŁŁśŁŁŁŮ ýďŘÀŁŁŁþć řc ŁŁŁČŁŁŁŮÂŁŁŁŮ ĆŁŁŁŞ ýćøďø ýŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ýĆŁŁŁţŁŁŁłď ÀŁŁŁłŚŁŁŁŞ ôěĎ ťŁŁŁĺŘ SSŁŁŁØŁŁŁÞŁŁŁõ ,üŁŁŁĺďŚŁŁŁê Ű cŚŁŁŁäĚŁŁŁÏŘ ěŘ ĆŁŁî üŁŁþŚŁŁûćÂŁŁŞďŚŁŁî .ÀŁŁĂŁŁî ýøďüŁŁŢ ÀŁŁîüŁŁœĄŁŁþ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ 9 ýĆŁŁÞŁŁČŁŁÞŁŁő ěŘ fŚŁŁÖŁŁČŁŁìć ÀŁŁþŚŁŁŞ ,ÛŁŁþÀŁŁśŁŁŮ SSŁŁþŘ ô cŚŁŁźŁŁœŘ ýùĄŁŁČŁŁł

.ÀŁĂŁţŁÆŁČŁœ ďŚŁðěŚŁĺ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ŚŁŞ ,ÀŁĂŁĂŁØŁœ ýøďüŁŢ ÂŁîÁŁóŘëĄŁê ýĆŁÞŁČŁÞŁő

üŁÚŁţŁÆŁŞ ćĄŁą éc ŘÂŁÏŘ ýŚŁûĆŁÆŁþĄŁœ ĆŁŞ ,"ěŚŁŞ ÃŁţŁœŘÂŁŢ’ ĆŁÜŁÞŁŻ ěŘ ,ŚŁûĆŁÆŁþĄŁœ ěŘ üŁÌŁãŁŞ Û cŁØŁł ŃŃŃ 1 ýďøŊćŚŁþ .ÀŁŁœĄŁŁłüŁŁõ ĆŁŁĂŁŁþÂŁŁì đĚŁŁÎŁŁňŘĆŁŁŞ ,ÀŁŁœÂŁŁČŁŁð ďŘÂŁŁì ÓŁŁóŚŁŁ£ŁŁõ ťc ŁŁúŁŁŻ ŚŁŁŞ ö cĄŁŁţŁŁõ ďć ĆŁŁî üŁŁŮďĄŁŁň ďć ŚŁŁûĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ .ćďŘć ùÀŁõŊ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ 9 ýĆŁÞŁČŁÞŁő ďć ďŚŁţŁêď SSŁþŘ đ c øÂŁÈŁõ .ÀŁĂŁĂŁî üŁœŚŁśŁČŁţŁÈŁŢ Řď ýěŚŁĺĆŁĂŁþÂŁì ÀŁþŚŁŞ ŚŁûćÂŁŞďŚŁî

.ťŁĺŘ

ýŘüŁŁĂŁŁÞŁŁő ý c ďŘÀŁŁþć řc ŁŁČŁŁŮÂŁŁŮ ĆŁŁî ÀŁŁþŊüŁŁõ ÇŁŁČŁŁŢ ýćďŘĄŁŁõ ,ĆŁŁţŁŁúŁŁŻøć ö cĄŁŁţŁŁõ ý c ďŘÁŁŁðÀŁŁî ďć ŃŃŃ 2 ýďøŊćŚŁŁþ ťc ŁŁŁúŁŁŁŻ ñ cÂŁŁŁţŁŁŁĂŁŁŁî ýŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ěŘ öŘĄŁŁŁŮüŁŁŁõ ŚŁŁŁûťŁŁŁóŚŁŁŁ§ SSŁŁŁþŘ ďć .ťŁŁŁÆŁŁŁČŁŁŁœ ŠĄŁŁŁÜŁŁŁÎŁŁŁõ ÀŁŁŁþŊüŁŁŁõ ťŁŁŁĺćĆŁŁŁŞ ŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ěŘ ĆŁŁŁî ÀŁœøďüŁõ ďŚŁîĆŁŞ SSŁţŁõ Ç c ŁþŚŁÞŁœ řc ŁČŁŮÂŁŮ ĳ c ŁČŁ¡ŁÊŁŮ ýŘÂŁŞ ÍŁÖŁê ŚŁûĆŁÆŁþĄŁœ SSŁþŘ .ćÂŁî ùćŚŁÔŁţŁĺŘ 1-5 Ç c Ł£ŁŞ 1 ñ cøÀŁŻ ýŘÂŁŁŞ .ÀŁŁœĄŁŁł ĆŁŁţŁŁêÂŁŁð ùÀŁŁþćŚŁŁœ (ŰŘďŚŁŁśŁŁä ýĄŁŁźŁŁţŁŁÆŁŁŻ ŚŁŁþ SSŁŁţŁŁõ ý c ěŚŁŁĺřŁŁŮÂŁŁõ À cŁŁĂŁŁœŚŁŁõ) ÂŁŁÚŁŁþć ýŚŁŁûľěŘćÂŁŁŢ ďć ÀŁŁþŚŁŁŞ ø

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ 9 ýĆŁÞŁČŁÞŁő ĆŁŞ ÂŁţŁÈŁČŁŞ Ű cŚŁäĚŁÏŘ

Š ťc ĺĄČŢ

ñŚŁÊŁŮŘ Ý cŁţŁþďĄŁÚŁóŘ

(üŁõŘÃŁóŘ)

ĆŁî üŁŮďĄŁň ďć ,ÀŁœÂŁČŁðüŁõ üŁÔŁÜŁţŁ£Łõ ñŚŁØŁłŘdćĄŁą ěŘ ÀŁãŁŞ ø ÛŁśŁì éc øÂŁ§ ĆŁŞ ĆŁţŁÆŁŞ ,üŁĺďŚŁê éc øÂŁ§ ĆŁî ŚŁźŁœŊ ěŘ Û cŁØŁł Š c ŚŁ£ŁţŁœŘ ø üŁĺďŚŁê Ű cŚŁäĚŁÏŘ Ç c ŁþŚŁÞŁœ ýŘÂŁŞ ,ÀŁĂŁûć ÇŁþŚŁÞŁœ üŁĺďŚŁê Íc Łą ŚŁŞ Řď ŚŁûĆŁÆŁþĄŁœ ÀŁĂŁûŘĄŁ£ŁŞ ŚŁûćÂŁŞďŚŁî Û cŁØŁł Â cŁČŁČŁçŁŮ Û cŁìŘÀŁ§ ,ÝŁţŁþďĄŁÚŁóŘ SSŁþŘ .ÀŁĂŁĂŁî ùćŚŁÔŁţŁĺŘ ťŁĺĄŁČŁŢ SSŁþŘ ďć ùÀŁł ÉŁ£ŁÈŁõ Ý cŁţŁþďĄŁÚŁóŘ ěŘ ÀŁþŚŁŞ ,řŁĺŚŁĂŁõ Ý cŁţŁþďĄŁÚŁóŘ ěŘ ,ćÂŁŞďŚŁî ĆŁŞ ĆŁţŁÆŁŞ ,ťŁĺŘ SSŁØŁÞŁõ üŁóø ,ÀŁĂŁîüŁõ ÉŁ£ŁÈŁõ üŁĺďŚŁê ö cĄŁţŁõ Ç c ŁþŚŁÞŁœ ýŘÂŁŞ Řď ěŚŁČŁœ ćcďĄŁõ

.(ÀŁĂŁûćüŁõ ÇŁþŚŁÞŁœ ÕŁČŁÜŁãŁţŁÆŁœ Íc Łą ŚŁŞ Řď SSŁţŁõ ĆŁî üŁþŚŁûćÂŁŞďŚŁî ďć fĚŁůŁõ) ćĄŁł ùćŚŁÔŁţŁĺŘ ÃŁČŁœ ýÂŁŮùÀŁČŁżŁČŁŢ ŚŁŁŁþ) ćĄŁŁŁł ôŚŁŁŁźŁŁŁœŘ ĆŁŁŁţŁŁŁúŁŁŁŻøć Ý cŁŁŁţŁŁŁþďĄŁŁŁÚŁŁŁóŘ ěŘ ÅŁŁŁŢ ,éŚŁŁŁÔŁŁŁł ýŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ SS cŁŁŁţŁŁŁêÂŁŁŁð ÂŁŁŁÑŁŁŁœďć ŚŁŁŁŞ ,ÀŁŁŁþŚŁŁŁŞ ñŚŁŁŁÊŁŁŁŮŘ Ý cŁŁŁţŁŁŁþďĄŁŁŁÚŁŁŁóŘ

.(ÀŁłŚŁŞ öŚŁÆŁØŁþ ćĄŁłüŁõ ôŚŁźŁœŘ ĆŁţŁúŁŻøć Ý cŁţŁþďĄŁÚŁóŘ ěŘ ÅŁŢ ÝŁţŁþďĄŁÚŁóŘ SSŁþŘ ĆŁî üŁţŁóŚŁ§ ŚŁŞ ÇŁČŁŻøÂŁą

ñŚŁÊŁŮŘ ýùćď 1-Š

Â cŁČŁŹŚbŁŮ ø ĆŁÆŁþĄŁœ Û cŁØŁł Â cŁČŁČŁçŁŮ ýùĄŁČŁł ,ŚŁûùćď SSŁþŘ .ćĄŁłüŁõ ýÀŁĂŁŞĆŁţŁĺć ñŚŁÊŁŮŘ ýŚŁûùćď ěŘ üŁØŁþ ďć ,ĆŁÆŁþĄŁœ ÂŁû

:ÀŁœÂŁþě đ c ÂŁł ĆŁŞ ŚŁûùćď SSŁþŘ .ÀŁĂŁĂŁîüŁõ ÉŁ£ŁÈŁõ ÂŁÚŁþć ýŚŁûĆŁÆŁþĄŁœ ÂŁŞ Řď öŊ

ŚŁŁŁŞ ."ćÂŁŁŁð Ű c’ ø ,"øŘø’ ,"ď c’ ,"ñŘć’ ,"ÓŁŁŁóŘ’ ,"Ŋ’ Û cŁŁŁČŁŁŁśŁŁŁì ěŘ üŁŁŁÜŁŁŁØŁŁŁłøć ýŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ :ÛŁŁŁňøťŁŁŁĺŘď •

.ÀŁœĄŁłüŁõ ÉŁ£ŁÈŁõ R éc ÂŁ§

ÉŁ£ŁÈŁõ D éc ÂŁ§ ŚŁŞ ."ćŚŁň’ ø ,"SSŁČŁĺ’ ,"ÝŁČŁŻ’ ,"Š c ’ Û cŁČŁśŁì ěŘ üŁÜŁØŁłďŚŁúŁİ ýŚŁûĆŁÆŁþĄŁœ :ÛŁňøøć •

.ÀŁœĄŁłüŁõ

SSŁŁþŘ Ű cøŚŁŁÔŁŁŮ ."ýěŚŁŁźŁŁõ ñ cŚŁŁÊŁŁŮŘ’ ø "üŁŁðÀŁŁČŁŁÈŁŁî’ Û cŁŁČŁŁśŁŁì ěŘ ,ñŚŁŁÊŁŁŮŘ ýùÀŁŁĂŁŁĂŁŁîćŚŁŁźŁŁþŘ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ :ÛŁŁňŘø • ÉŁŁŁ£ŁŁŁÈŁŁŁõ C éc ÂŁŁŁ§ ŚŁŁŁŞ .ÀŁŁŁĂŁŁŁûćüŁŁŁÞŁŁŁœ ÛŁŁŁØŁŁŁł Â cŁŁŁČŁŁŁČŁŁŁçŁŁŁŮ ĆŁŁŁî ťŁŁŁĺŘ SSŁŁŁþŘ "ÛŁŁŁňøøć’ ýŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ŚŁŁŁŞ ŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ

.ÀŁœĄŁłüŁõ Â cŁŁČŁŁè ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ýĆŁŁČŁŁÜŁŁî ø "ýěŚŁŁźŁŁõ ýĆŁŁÜŁŁňŚŁŁê’ Û cŁŁõŚŁŁł ,ñŚŁŁÊŁŁŮŘ ýùÀŁŁĂŁŁĂŁŁîâŁŁÎŁŁì ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ :ÛŁŁňŚŁŁê • Ý cŁŁŁŐĚŁŁŁä ,ôŚŁŁŁìďŘ ,ŚŁŁŁûĆŁŁŁÜŁŁŁňŚŁŁŁê ,"ùÃŁŁŁÞŁŁŁû’ Û cŁŁŁČŁŁŁśŁŁŁì ěŘ ,ÀŁŁŁœÂŁŁŁČŁŁŁðüŁŁŁÞŁŁŁœ ďŘÂŁŁŁì ëĄŁŁŁê ý c ÀŁŁŁĂŁŁŁŞĆŁŁŁţŁŁŁĺć ďć ĆŁŁŁî ùÀŁŁŁœĄŁŁŁłřŁŁŁČŁŁŁîÂŁŁŁŮ

.ÀŁœĄŁłüŁõ ÉŁ£ŁÈŁõ U éc ÂŁ§ ŚŁŞ .ùÂŁČŁè ø üŁœŚŁœĄŁþ ,SSŁČŁŮĎ ýŚŁûÍŁą éc øÂŁ§ ø ,ýďŘÁŁðĆŁÎŁÖŁœ

Û cŁŁČŁŁśŁŁì ěŘ ,üŁŁóÂŁŁţŁŁĂŁŁî ø ùÀŁŁœĄŁŁłřŁŁČŁŁîÂŁŁŮ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ Û cŁŁõŚŁŁł ,ñŚŁŁÊŁŁŮŘ ĆŁŁŞ ťŁŁśŁŁÆŁŁœ éŚŁŁÔŁŁł ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ :éŚŁŁÔŁŁł • T éc ÂŁŁ§ ŚŁŁŞ ."šŁŁİĆŁŁŞťŁŁĺŘď ýĆŁŁœŚŁŁÈŁŁœ’ ø ,"ùďĄŁŁÊŁŁÖŁŁõ Ó c ŁŁóŘ’ ,"ÀŁŁþÀŁŁÈŁŁŮ’ ,"öĄŁŁØŁŁĺ’ ,"ÂŁŁŞděøć’ ,"ÂŁŁŞdě’

.ÀŁœĄŁłüŁõ ÉŁ£ŁÈŁõ

đ c ĚŁŁŁÎŁŁŁňŘ ěŘ ø ;"ÛŁŁŁňŘø’ ø "ÛŁŁŁňøøć’ ýŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ýŘÂŁŁŁŞ "šŁŁŁİĆŁŁŁŞÛŁŁŁÊŁŁŁţŁŁŁõ’ đ c ĚŁŁŁÎŁŁŁňŘ ěŘ ,ťŁŁŁĺĄŁŁŁČŁŁŁŢ SSŁŁŁþŘ ďć

.ćĄŁłüŁõ ùćŚŁÔŁţŁĺŘ "ÛŁňŘø’ ø ,"ÛŁňøøć’ ,"ÛŁňøťŁĺŘď’ ýŚŁûĆŁÆŁþĄŁœ ýŘÂŁŞ "ťŁĺŘďĆŁŞÛŁÊŁţŁõ’

ĆŁŁŁî ,ÀŁŁŁîüŁŁŁœĄŁŁŁþ ý c ŘùćŘć ýŚŁŁŁûùÀŁŁŁœøÂŁŁŁŢ ěŘ ArabicShaping.txt ýùÀŁŁŁœøÂŁŁŁŢ ğ c ŚŁŁŁĺŘ ÂŁŁŁŞ ÀŁŁŁþŚŁŁŁŞ ,ŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ýùćď

ü c ŁţŁœÂŁţŁĂŁþŘ ü c ŁœŚŁÈŁœ ďć öŊ ýĆŁ£ŁÆŁœ SSŁþÂŁąŊ

          http://www.unicode.org/Public/UNIDATA/ArabicShaping.txt

.ćĄŁł SSŁČŁČŁãŁŮ ťŁĺŘ ğÂŁţŁĺć ďć

ŚŁŁûĆŁŁÆŁŁþĄŁŁœ Û cŁŁØŁŁł Â cŁŁČŁŁČŁŁçŁŁŮ ýŘÂŁŁŞ "ýěŚŁŁźŁŁõ ñ cŚŁŁÊŁŁŮŘ’ ø "ýěŚŁŁźŁŁõ ýĆŁŁÜŁŁňŚŁŁê’ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ŃŃŃ ýďøŊćŚŁŁþ ŚŁþ ,("ŚŁûĆŁœŚŁą’ ýĆŁÞŁÜŁî ďć fĚŁůŁõ) ďøŚŁźŁõ éc ÂŁ§øć ñ cŚŁÊŁŮŘ ěŘ ťŁãŁœŚŁÞŁõ ýŘÂŁŞ ĆŁÆŁþĄŁœ øć SSŁþŘ ěŘ .ÀŁœøďüŁõ ďŚŁîĆŁŞ ("üŁŁÆŁŁÞŁŁł ý c ÂŁŁźŁŁû’ Ó c ŁŁÔŁŁ£ŁŁõ ö cŘĄŁŁĂŁŁäĆŁŁŞ ,".ľ .ŁŁû’ ďć fĚŁŁůŁŁõ) éøÂŁŁ§ ñ cĄŁŁÞŁŁãŁŁõ Û cŁŁØŁŁł ěŘ ÂŁŁČŁŁè üŁŁÜŁŁØŁŁł Š c ŚŁŁ£ŁŁţŁŁœŘ

.ćĄŁłüŁõ ùćŚŁÔŁţŁĺŘ

ÝŁţŁþďĄŁÚŁóŘ 2-Š

,"ñøŘ’ đĚŁÎŁňŘ ďć ŚŁûÛŁØŁł SSŁþŘ .ćĄŁłüŁõ SSŁČŁČŁãŁŮ ÛŁØŁł ďŚŁúŁİ ŚŁŮ ,öŚŁÈŁóŚŁÊŁŮŘ ýùćď ĆŁŞ ĆŁţŁÆŁŞ ,ŚŁûĆŁÆŁþĄŁœ ýŘÂŁŞ Řď "ŚŁúŁĂŁŮ’ ø "ÂŁąŊ’ Û cŁØŁł øć ÍŁÖŁê "ÛŁňøťŁĺŘď’ ýŚŁûĆŁÆŁþĄŁœ .ÀŁœĄŁłüŁõ ùÀŁČŁõŚŁœ "ŚŁúŁĂŁŮ’ ø ,"ÂŁąŊ’ ,"ÍŁĺø’

.ÀŁœÂŁČŁðüŁõ

:ťŁĺŘ ÂŁþě đ c ÂŁł ĆŁŞ ñŚŁÊŁŮŘ Ý cŁţŁþďĄŁÚŁóŘ

,ÀŁŁãŁŁŞ ĆŁŁŞ SSŁŁþŘ ěŘ SSŁŁþŘÂŁŁŞŚŁŁĂŁŁŞ) .ÀŁŁĂŁŁûćüŁŁÞŁŁœ ÂŁŁČŁŁČŁŁçŁŁŮ Řď ĆŁŁþŚŁŁŢ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ü c ŁŁóŚŁŁÊŁŁŮŘ ď cŚŁŁţŁŁêď "éŚŁŁÔŁŁł’ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ .1 ďć ďĄŁÏSSŁČŁÞŁû ;ćĄŁŞ ÀŁûŘĄŁą ťŁĺŘď ťc ŁÞŁĺ éc ŚŁÔŁł Â cŁČŁè ýĆŁÆŁþĄŁœ SSŁČŁóøŘ ,ťŁĺŘď ťc ŁÞŁĺ ýĆŁÆŁþĄŁœ ěŘ ďĄŁÑŁĂŁõ

(.šŁİ ťc ŁÞŁĺ ýĆŁÆŁþĄŁœ ćcďĄŁõ Û cŁŁŁØŁŁŁł ĆŁŁŁŞ ,ÀŁŁŁłŚŁŁŁŞ "šŁŁŁİĆŁŁŁŞÛŁŁŁÊŁŁŁţŁŁŁõ’ ÇŁŁŁţŁŁŁĺŘď ťc ŁŁŁÞŁŁŁĺ ýĆŁŁŁÆŁŁŁþĄŁŁŁœ ø ,ÀŁŁŁłŚŁŁŁŞ "ÛŁŁŁňøťŁŁŁĺŘď’ ýŘĆŁŁŁÆŁŁŁþĄŁŁŁœ ÂŁŁŁðŘ .2

.ÀŁþŊüŁõ ďć "ÂŁąŊ’

ťc ŁŁÞŁŁĺ ýĆŁŁÆŁŁþĄŁŁœ ø ,ÀŁŁłŚŁŁŞ "šŁŁİĆŁŁŞÛŁŁÊŁŁţŁŁõ’ ÇŁŁţŁŁĺŘď ťc ŁŁÞŁŁĺ ýĆŁŁÆŁŁþĄŁŁœ ,ÀŁŁłŚŁŁŞ "ÛŁŁňøøć’ ýŘĆŁŁÆŁŁþĄŁŁœ ÂŁŁðŘ .3

.ÀŁþŊüŁõ ďć "ÍŁĺø’ Û cŁØŁł ĆŁŞ ,ÀŁłŚŁŞ "ťŁĺŘďĆŁŞÛŁÊŁţŁõ’ ÇŁşŁİ

ťc ŁŁÞŁŁĺ ýĆŁŁÆŁŁþĄŁŁœ ø ,ÀŁŁłŚŁŁŞ "šŁŁİĆŁŁŞÛŁŁÊŁŁţŁŁõ’ ÇŁŁţŁŁĺŘď ťc ŁŁÞŁŁĺ ýĆŁŁÆŁŁþĄŁŁœ ,ÀŁŁłŚŁŁŞ "ÛŁŁňøøć’ ýŘĆŁŁÆŁŁþĄŁŁœ ÂŁŁðŘ .4

.ÀŁþŊüŁõ ďć "ÂŁąŊ’ Û cŁØŁł ĆŁŞ ,ÀŁłŚŁśŁœ "ťŁĺŘďĆŁŞÛŁÊŁţŁõ’ ÇŁşŁİ

ťc ŁÞŁĺ ýĆŁÆŁþĄŁœ ø ,ÀŁłŚŁśŁœ "šŁİĆŁŞÛŁÊŁţŁõ’ ÇŁţŁĺŘď ťc ŁÞŁĺ ýĆŁÆŁþĄŁœ ,ÀŁłŚŁŞ "ÛŁňøøć’ ýŘĆŁÆŁþĄŁœ ÂŁðŘ .5

.ÀŁþŊüŁõ ďć "ñøŘ’ Û cŁØŁł ĆŁŞ ,ÀŁłŚŁŞ "ťŁĺŘďĆŁŞÛŁÊŁţŁõ’ ÇŁşŁİ

.ÀŁþŊüŁõ ďć "ŚŁúŁĂŁŮ’ Û cŁØŁł ĆŁŞ ĆŁÆŁþĄŁœ ,ÀŁĂŁłŚŁśŁœ ďŘÂŁìÂŁŞ ëĄŁê ýŚŁûťŁóŚŁ§ ěŘ ŒŁþžŁČŁû ĆŁî üŁŮďĄŁň ďć .6

,ÀŁŁŁłŚŁŁŁŞ ćĄŁŁŁą À cŁŁŁĂŁŁŁŞ ŚŁŁŁþ ÂŁŁŁÎŁŁŁĺ "éŚŁŁŁÔŁŁŁł’Â cŁŁŁČŁŁŁè ýĆŁŁŁÆŁŁŁþĄŁŁŁœ SSŁŁŁČŁŁŁóøŘ ýŘĆŁŁŁÆŁŁŁþĄŁŁŁœ ĆŁŁŁî üŁŁŁŮďĄŁŁŁň ďć ŃŃŃ 1 ýďøŊćŚŁŁŁþ ýĆŁŁÆŁŁþĄŁŁœ SSŁŁþÂŁŁąŊ ýŘĆŁŁÆŁŁþĄŁŁœ ĆŁŁî üŁŁŮďĄŁŁň ďć ,ďĄŁŁÏSSŁŁČŁŁÞŁŁû .ćĄŁŁłüŁŁõ ŋÂŁŁê "ÛŁŁňŚŁŁê’ ÇŁŁţŁŁĺŘď ťc ŁŁÞŁŁĺ ýĆŁŁÆŁŁþĄŁŁœ

.ćĄŁłüŁõ ŋÂŁê "ÛŁňŚŁê’ ÇŁşŁİ ťc ŁÞŁĺ ýĆŁÆŁþĄŁœ ,ÀŁłŚŁŞ ćĄŁą À cŁĂŁŞ ŚŁþ ÂŁÎŁĺ "éŚŁÔŁł’Â cŁČŁè ýŚŁŁûĆŁŁÆŁŁþĄŁŁœ ,ćĄŁŁłüŁŁõ ôŚŁŁźŁŁœŘ ĆŁŁţŁŁúŁŁŻøć Ý cŁŁţŁŁþďĄŁŁÚŁŁóŘ ěŘ ÅŁŁŢ ÝŁŁţŁŁþďĄŁŁÚŁŁóŘ SSŁŁþŘ ĆŁŁî ŚŁŁŻöŊ ěŘ ŃŃŃ 2 ýďøŊćŚŁŁþ

.ÀŁœĄŁłüŁõ SSŁČŁČŁãŁŮ ýďŘÀŁþć řc ŁČŁŮÂŁŮ ğ c ŚŁĺŘ ÂŁŞ šŁİ ø ťŁĺŘď ťc ŁÞŁĺ

ñŚŁÊŁŮŘ ù cøÂŁð 3-Š

ñŚŁÊŁŮŘ ýŚŁûùøÂŁð ďć ýÂŁûŚŁÒ Û cŁØŁł ĆŁŞ ĆŁţŁÆŁŞ ,ÀŁœÂŁČŁðüŁõ ÓŁÜŁţŁ£Łõ ýŚŁûÛŁØŁł ĆŁî ýŘüŁĺďŚŁê éc øÂŁ§ ěŘ ŒŁþ ÂŁû ýŚŁŁŁûùÀŁŁŁœøÂŁŁŁŢ ěŘ ArabicShaping.txt ýùÀŁŁŁœøÂŁŁŁŢ ğ c ŚŁŁŁĺŘ ÂŁŁŁŞ ÀŁŁŁþŚŁŁŁŞ ÃŁŁŁČŁŁŁœ ŚŁŁŁûùøÂŁŁŁð SSŁŁŁþŘ .ÀŁŁŁœĄŁŁŁłüŁŁŁõ ýÀŁŁŁĂŁŁŁŞĆŁŁŁţŁŁŁĺć

ü c ŁţŁœÂŁţŁĂŁþŘ ü c ŁœŚŁÈŁœ ďć öŊ ýĆŁ£ŁÆŁœ SSŁþÂŁąŊ ĆŁî ,ÀŁîüŁœĄŁþ ý c ŘùćŘć

          http://www.unicode.org/Public/UNIDATA/ArabicShaping.txt

.ćĄŁł SSŁČŁČŁãŁŮ ťŁĺŘ ğÂŁţŁĺć ďć

ďć ĆŁŁŁŁî ýÂŁŁŁŁþÁŁŁŁŁŢÛŁŁŁŁØŁŁŁŁł éc øÂŁŁŁŁ§ ,ćďŘÀŁŁŁŁœŚŁŁŁŁţŁŁŁŁĺŘ SSŁŁŁŁþŘ SS cŁŁŁŁþøÀŁŁŁŁŮ ö cŚŁŁŁŁõě ďć ğďťŁŁŁŁĺć ďć ýĆŁŁŁŁ£ŁŁŁŁÆŁŁŁŁœ SSŁŁŁŁþÂŁŁŁŁąŊ ğ c ŚŁŁŁŁĺŘ ÂŁŁŁŁŞ :ÀŁœÂŁČŁðüŁõ ďŘÂŁì ŚŁûùøÂŁð SSŁþŘ ďć ÀŁœŘùÀŁõŊ 6-5 ø 5-5 ýŚŁûÇŁ£ŁŞ

éøÂŁ§ ñ cŚŁÊŁŮŘ ýŚŁûùøÂŁð ťc ŁĺÂŁúŁê ŃŃŃ 9 ñ cøÀŁŻ

ŚŁûĆŁÆŁþĄŁœ  ñŚŁÊŁŮŘ ýùćď ùøÂŁð SS cŁČŁŮĎ ô cŚŁœ ùøÂŁð ô cŚŁœ

,"ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ’ ,"ÓŁóŘ’ ,"Ŋ’ÛŁňøťŁĺŘď ALEF ÓŁóŘ "ÛŁňø Ó c ŁóŘ’ ,"SSŁČŁþŚŁŢ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ’

"Ÿ c’ ø ,"Ű c’ ,"Ť c’ ,"Š c ’  ÛŁňøøć  BEH

Š c

"ă c ’ ø ,"đ c ’ ,"Ĳ c ’ ,"ÝŁČŁŻ’  ÛŁňøøć  HAH

đ c

"ñŘč’ ø ,"ñŘć’ ÛŁňøťŁĺŘď DAL  ñŘć
"ęc’ ø ,"ěc’ ,"ď c’  ÛŁňøťŁĺŘď REH  ď c
"SSŁČŁł’ ø ,"SSŁČŁĺ’ ÛŁňøøć  SEEN  SSŁČŁĺ
"ćŚŁő’ ø ,"ćŚŁň’  ÛŁňøøć  SAD  ćŚŁň
"ŚŁÒ’ ø ,"ŚŁÏ’  ÛŁňøøć  TAH  ŚŁÏ
"SSŁČŁè’ ø ,"SSŁČŁä’ ÛŁňøøć  AIN  SSŁČŁä
"éc ’  ÛŁňøøć  FEH  éc
"éŚŁì’  ÛŁňøøć  QAF  éŚŁì
"üŁŞÂŁä éc ŚŁî’ ÛŁňøøć  KAF  üŁŞÂŁä éc ŚŁî
"éŚŁð’ ø ,"éŚŁî’  ÛŁňøøć  GAF  éŚŁð
"ôĎ’  ÛŁňøøć  LAM  ôĎ
"ÝŁČŁõ’ ÛŁňøøć  MEEM  ÝŁČŁõ
"öĄŁœ’ ÛŁňøøć  NOON  öĄŁœ

"ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ øŘø’ ø ,"øŘø’

ÛŁňøťŁĺŘď WAW øŘø

"Łû c’ ÛŁňøøć HEH Ł û c
"ćÂŁð Ű c’ ÛŁňøťŁĺŘď TEH MARBUTA  ćÂŁð Ű c

,"ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c ’ ,"ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ý c ’ ,"ý c ’ ÛŁňøøć YEH

ý c

"ĆŁÎŁÖŁœüŁŞ ü c ŁŞÂŁä ý c ’ ø

ćĄŁŁŻĄŁŁõ Ű cŚŁŁäĚŁŁÏŘ ĆŁŁŞ ÀŁŁĂŁŁÔŁŁÒĄŁŁõ ŚŁŁûćÂŁŁŞďŚŁŁî .ćďŘć üŁŁŮŚŁŁäĚŁŁÏŘ ýĆŁŁśŁŁĂŁŁŻ ÍŁŁÖŁŁê ëĄŁŁê ťc ŁŁĺÂŁŁúŁŁê ŃŃŃ ýďøŊćŚŁŁþ .ÀŁĂŁĂŁî ĆŁãŁŻŘÂŁõ ArabicShaping.txt ýùÀŁœøÂŁŢ ďć

ŚŁûďĄŁŮŚŁÚŁČŁó 4-Š

fĚŁŁůŁŁõ .ÀŁŁœěŚŁŁÆŁŁŞ ďĄŁŁŮŚŁŁÚŁŁČŁŁó ô cŚŁŁœ ĆŁŁŞ ýŘüŁŁêÂŁŁ§ÀŁŁĂŁŁİ ñ cŚŁŁØŁŁłŘd,ùćŚŁŁÔŁŁţŁŁĺŘ ćcďĄŁŁõ Ý cŁŁÜŁŁì ĆŁŁŞ ĆŁŁţŁŁÆŁŁŞ ÀŁŁĂŁŁœŘĄŁŁŮüŁŁõ üŁŁĺďŚŁŁê éc øÂŁŁ§

.ÀŁþŚŁČŁŞ " ’ Û cŁØŁłĆŁŞ ÀŁœŘĄŁŮüŁõ "ý c ’ ø "éc ’ řc ŁČŁîÂŁŮ ø ,"Ď’ Û cŁØŁłĆŁŞ ÀŁœŘĄŁŮüŁõ "ÓŁóŘ’ ø "ôĎ’ řŁČŁîÂŁŮ ĆŁî ÀŁĂŁţŁÆŁû üŁþŚŁûďĄŁŮŚŁÚŁČŁó ,ýďŚŁśŁŻŘ ýŚŁûďĄŁŮŚŁÚŁČŁó .ÀŁœŘýďŚŁśŁŻŘ ÂŁÚŁþć üŁÌŁãŁŞ ø ýďŚŁČŁţŁąŘ ŚŁûďĄŁŮŚŁÚŁČŁó ěŘ üŁÌŁãŁŞ .ÀŁŁŁłŚŁŁŁŞ (ALEF) "ÓŁŁŁóŘ’ ñ cŚŁŁŁÊŁŁŁŮŘ ù cøÂŁŁŁð ěŘ öŚŁŁŁÈŁŁŁõøć éc ÂŁŁŁ§ ø (LAM) "ôĎ’ ñ cŚŁŁŁÊŁŁŁŮŘ ù cøÂŁŁŁð ěŘ öŚŁŁŁÈŁŁŁóøŘ éc ÂŁŁŁ§ ďĄŁŁŮŚŁŁÚŁŁČŁŁó Û cŁŁØŁŁł ĆŁŁŞ ťŁŁĺŘ SSŁŁØŁŁÞŁŁõ üŁŁÈŁŁþŚŁŁÞŁŁœ Ý cŁŁÜŁŁì ĆŁŁŞ ĆŁŁţŁŁÆŁŁŞ ĆŁŁî ÀŁŁĂŁŁţŁŁÆŁŁû ÂŁŁÚŁŁþć ýŚŁŁûďĄŁŁŮŚŁŁÚŁŁČŁŁó ,ýďŚŁŁČŁŁţŁŁąŘ ýŚŁŁûďĄŁŁŮŚŁŁÚŁŁČŁŁó ,ÀŁĂŁûć ÇŁþŚŁÞŁœ ďĄŁŮŚŁÚŁČŁó Û cŁØŁłĆŁŞ Řď ýďŚŁśŁŻŘ ýŚŁûďĄŁŮŚŁÚŁČŁó ,öĄŁţŁõ Ç c ŁþŚŁÞŁœ ďć ÀŁĂŁÔŁÒĄŁõ ŚŁûćÂŁŞďŚŁî .ÀŁĂŁŞŚŁþ ÇŁþŚŁÞŁœ ýŘùÄŁŁþø ýŚŁŁûťŁŁþćøÀŁŁ¡ŁŁõ ÇŁŁþŚŁŁÞŁŁœ ù cŚŁŁÚŁŁţŁŁĺć ŚŁŁþ ,ÀŁŁłŚŁŁŞ ÂŁŁÑŁŁœ ćcďĄŁŁõ ńŚŁŁą ý c ÂŁŁÊŁŁŞ ýŚŁŁûùĄŁŁÜŁŁŻ ĆŁŁî ýćďŘĄŁŁõ ďć ÂŁŁÚŁŁõ

.ÀŁłŚŁŞ ĆŁţŁłŘć

:ćĄŁłüŁõ ùćŚŁÔŁţŁĺŘ ÂŁþě Ý cŁţŁþďĄŁÚŁóŘ ěŘ ,ŚŁûďĄŁŮŚŁÚŁČŁó SSŁþŘ ñ cŚŁÞŁäŘ cýŘÂŁŞ

.ÀŁĂŁûćüŁÞŁœ ÂŁČŁČŁçŁŮ Řď ĆŁþŚŁŢ ýŚŁûĆŁÆŁþĄŁœ ý c ďĄŁŮŚŁÚŁČŁó ď cŚŁţŁêď "éŚŁÔŁł’ ýŚŁûĆŁÆŁþĄŁœ .1

ø ,"ÍŁŁĺø’ Û cŁŁØŁŁłĆŁŁŞ ø "ôĎ’ ù cøÂŁŁð ďć ÇŁŁţŁŁĺŘď ťŁŁÞŁŁĺ ýĆŁŁÆŁŁþĄŁŁœ ĆŁŁî ŚŁŁûĆŁŁÆŁŁþĄŁŁœ ěŘ üŁŁþŚŁŁŮøć ýĆŁŁóŚŁŁśŁŁœć ÂŁŁû .2 Řď "ÓŁóŘôĎ’ ýĆŁţŁĺć ěŘ ýďĄŁŮŚŁÚŁČŁó ,ÀŁłŚŁŞ "ÂŁąŊ’ Û cŁØŁłĆŁŞ ø "ÓŁóŘ’ ù cøÂŁð ďć ÇŁşŁİ ťc ŁÞŁĺ ýĆŁÆŁþĄŁœ

.ÀŁûćüŁõ ÛŁČŁØŁÈŁŮ "ÂŁąŊ’ Û cŁØŁłĆŁŞ

ø ,"ñøŘ’ Û cŁŁŁØŁŁŁłĆŁŁŁŞ ø "ôĎ’ ù cøÂŁŁŁð ďć ÇŁŁŁţŁŁŁĺŘď ťŁŁŁÞŁŁŁĺ ýĆŁŁŁÆŁŁŁþĄŁŁŁœ ĆŁŁŁî ŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ěŘ üŁŁŁþŚŁŁŁŮøć ýĆŁŁŁóŚŁŁŁśŁŁŁœć ÂŁŁŁû .3 Řď "ÓŁóŘôĎ’ ýĆŁţŁĺć ěŘ ýďĄŁŮŚŁÚŁČŁó ,ÀŁłŚŁŞ "ÂŁąŊ’ Û cŁØŁłĆŁŞ ø "ÓŁóŘ’ ù cøÂŁð ďć ÇŁşŁİ ťc ŁÞŁĺ ýĆŁÆŁþĄŁœ

.ÀŁûćüŁõ ÛŁČŁØŁÈŁŮ "ŚŁúŁĂŁŮ’ Û cŁØŁłĆŁŞ

,ÀŁœĄŁłüŁÞŁœ ďĄŁŮŚŁÚŁČŁó üŁóø ÀŁœĄŁłüŁõ ÛŁÊŁţŁõ ÝŁû ĆŁŞ ŋÂŁÔŁÈŁČŁŢ ď cĄŁÏĆŁŞ ĆŁî ĆŁÆŁþĄŁœ ÀŁĂŁİ ŚŁþ øć SS cŁČŁŞ ùŚŁð ÂŁû .4 ý c ďĄŁŁŁŁŮŚŁŁŁŁÚŁŁŁŁČŁŁŁŁó Û cŁŁŁŁØŁŁŁŁł ö cćĄŁŁŁŁŞ ćĄŁŁŁŁŻĄŁŁŁŁõ Ű cďĄŁŁŁŁň ďć ,ćÂŁŁŁŁČŁŁŁŁð ďŘÂŁŁŁŁì "ýěŚŁŁŁŁźŁŁŁŁõ ñ cŚŁŁŁŁÊŁŁŁŁŮŘ’ ýĆŁŁŁŁÆŁŁŁŁþĄŁŁŁŁœ ÀŁŁŁŁĂŁŁŁŁİ ŚŁŁŁŁþ ŒŁŁŁŁþ

.ćĄŁł ùćŚŁÔŁţŁĺŘ ýďĄŁŮŚŁÚŁČŁó Û cŁØŁł ěŘ ÀŁþŚŁŞ ,ÇŁþŚŁÞŁœ ýŘÂŁŞ ùćŚŁÔŁţŁĺŘ ćcďĄŁõ Ý cŁÜŁì ďć ŚŁûĆŁÆŁþĄŁœ

ŚŁûĆŁÆŁþĄŁœ ý c ćŚŁä Û cŁØŁł ÀŁþŚŁŞ ,ćÂŁČŁð ďŘÂŁì "ýěŚŁźŁõ ýĆŁÜŁňŚŁê’ ýĆŁÆŁþĄŁœ ,ĆŁÆŁþĄŁœ ÀŁĂŁİ ŚŁþ øć SS cŁČŁŞ ùŚŁð ÂŁû .5 "ýěŚŁŁźŁŁõ ñ cŚŁŁÊŁŁŮŘ ,ýěŚŁŁźŁŁõ ýĆŁŁÜŁŁňŚŁŁê ,ýěŚŁŁźŁŁõ ñ cŚŁŁÊŁŁŮŘ’ ýĆŁŁóŚŁŁśŁŁœć ÂŁŁðŘ fĚŁŁůŁŁõ .ćÂŁŁČŁŁð ďŘÂŁŁì ùćŚŁŁÔŁŁţŁŁĺŘ ćďĄŁŁõ ÇŁþŚŁÞŁœ "ŚŁó’ Û cŁØŁł ĆŁŞ ÀŁþŚŁŞ ĆŁØŁÜŁŞ ,ÀŁœĄŁł ďĄŁŮŚŁÚŁČŁó ÀŁþŚŁśŁœ éÂŁ§ øć SSŁþŘ ,ćÂŁČŁð ďŘÂŁì "ÓŁóŘ’ ø "ôĎ’ SS cŁČŁŞ

.ÀŁĂŁŞŚŁþ

Ť ťc ĺĄČŢ

ŚŁûùćŘć ñćŚŁśŁŮ ýŚŁûřŁóŚŁì

(üŁõŘÃŁóŘ)

ćcďŘÀŁŁœŚŁŁţŁŁĺŘ üŁŁóø .ćÂŁŁî ùćŚŁŁÔŁŁţŁŁĺŘ ŰŚŁŁäĚŁŁÏŘ ñ cćŚŁŁśŁŁŮ ýŘÂŁŁŞ ýćÀŁŁãŁŁţŁŁõ ýŚŁŁûřŁŁóŚŁŁì ěŘ öŘĄŁŁŮüŁŁõ ÀŁŁîüŁŁœĄŁŁþ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ ďć ,ÕŁŁţŁŁÈŁŁõ ýŚŁŁûřŁŁóŚŁŁì .ćěŘćÂŁŁŢüŁŁõ ,UTF-32 ø ,UTF-16 ,UTF-8 üŁŁĂŁŁãŁŁþ ,üŁŁÜŁŁňŘ ýŚŁŁûřŁŁóŚŁŁì ĆŁŁŞ ÍŁŁÖŁŁê ÂŁŁőŚŁŁ§

.ÀŁœĄŁłüŁõ ŋÂŁê (UTF-16 ñŚŁůŁõ SSŁþŘ ďć) öŚŁłÂŁČŁÑŁœ üŁÜŁňŘ řc ŁóŚŁì ěŘ üŁňŚŁą ťc ŁóŚŁ§ ,UTF-16LE fĚŁůŁõ ďć ÀŁŁŁŁĂŁŁŁŁÔŁŁŁŁÒĄŁŁŁŁõ ,ÀŁŁŁŁĂŁŁŁŁĂŁŁŁŁîüŁŁŁŁõ ñĄŁŁŁŁśŁŁŁŁì ýćøďø ŚŁŁŁŁþ ,ÀŁŁŁŁĂŁŁŁŁûćüŁŁŁŁõ üŁŁŁŁŻøÂŁŁŁŁą ýÀŁŁŁŁîüŁŁŁŁœĄŁŁŁŁþ ýŚŁŁŁŁûřŁŁŁŁóŚŁŁŁŁì ďć ĆŁŁŁŁî üŁŁŁŁþŚŁŁŁŁûćÂŁŁŁŁŞďŚŁŁŁŁî ø ,RFC 2781 ,RFC 2279 ěŘ řŁŁČŁŁŮÂŁŁŮĆŁŁŞ ,UTF-32 ø ,UTF-16 ,UTF-8 ýŚŁŁûřŁŁóŚŁŁì ěŘ ùćŚŁŁÔŁŁţŁŁĺŘ Ű cďĄŁŁň ùćŚŁŁŁÔŁŁŁţŁŁŁĺŘ UTF-8 řc ŁŁŁóŚŁŁŁì ěŘ ŚŁŁŁûćÂŁŁŁŞďŚŁŁŁî ĆŁŁŁî üŁŁŁŮďĄŁŁŁň ďć .ÀŁŁŁĂŁŁŁĂŁŁŁî ýøďüŁŁŁŢ ÀŁŁŁîüŁŁŁœĄŁŁŁþ ćcďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ 19 ýĆŁŁŁÞŁŁŁČŁŁŁÞŁŁŁő üŁŮďĄŁň ďć ťŁĺŘ ÂŁţŁúŁŞ üŁóø ,ÀŁĂŁĂŁØŁœ ÀŁČŁóĄŁŮ üŁŻøÂŁą ýŘÀŁţŁŞŘ ďć Řď U+FEFF ýĆŁÆŁþĄŁœ ťŁĺŘ ÂŁţŁúŁŞ ,ÀŁĂŁĂŁîüŁõ ďć ùÀŁŁĂŁŁĂŁŁîÉŁŁ£ŁŁÈŁŁõ ťc ŁŁõĚŁŁä ö cŘĄŁŁĂŁŁäĆŁŁŞ Řď öŊ ,ÀŁŁþŚŁŁČŁŁŞ UTF-8 řc ŁŁóŚŁŁì ďć ýŚŁŁûýćøďø ýŘÀŁŁţŁŁŞŘ ďć ĆŁŁÆŁŁþĄŁŁœ SSŁŁþŘ ĆŁŁî

.ÀŁĂŁĂŁØŁœ ÇŁłěŘćÂŁŢ ø ,ĆŁţŁêÂŁð ÂŁÑŁœ

15 ýĆŁŁÞŁŁČŁŁÞŁŁő ďć ùÀŁŁł ÉŁŁ£ŁŁÈŁŁõ ñ cŚŁŁõÂŁŁœ ýŚŁŁûŰďĄŁŁň ěŘ üŁŁØŁŁþ ťŁŁĺŘ ÂŁŁţŁŁúŁŁŞ ŚŁŁûćÂŁŁŞďŚŁŁî ŃŃŃ ýďøŊćŚŁŁþ

ÂŁţŁÈŁČŁŞ Ű cŚŁäĚŁÏŘ ýŘÂŁŞ) .ÀŁĂŁĂŁî ÀŁČŁóĄŁŮ řŁóŚŁì öŊ ďć Řď ćĄŁą ýŚŁûüŁŻøÂŁą ø ùćÂŁî ŠŚŁ£ŁţŁœŘ Řď ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ (.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ Ÿ ťc ŁĺĄŁČŁŢ ĆŁŞ

Ű ťc ĺĄČŢ

ýÀŁĂŁŞéŘÂŁðŘďŚŁŢ ø ýÀŁĂŁŞÂŁÎŁĺ

(üŁõŘÃŁóŘ)

ćcďŘÀŁœŚŁţŁĺŘ 13 ýĆŁÞŁČŁÞŁő ýŚŁûĆŁČŁňĄŁŮ ğ c ŚŁĺŘ ÂŁŞ Řď ŚŁûÀŁĂŁŞ ø ŚŁûÂŁÎŁĺ ýùÀŁĂŁĂŁîŘÀŁŻ ýŚŁûĆŁÆŁþĄŁœ ÀŁĂŁÔŁÒĄŁõ ŚŁûćÂŁŞďŚŁî

.ÀŁĂŁĂŁî ÂŁČŁÆŁÔŁŮ ÀŁîüŁœĄŁþ

ďć üŁĂŁţŁõ ÀŁûŘĄŁ£ŁŞ ýćÂŁŞďŚŁî ĆŁî üŁŮďĄŁň ďć ,ÂŁîÁŁóŘëĄŁê ýĆŁÞŁČŁÞŁő ďć ùÀŁłÉŁ£ŁÈŁõ ýŚŁûťŁþćøÀŁ¡Łõ ÂŁŞ ùøĚŁä "ŚŁûÀŁĂŁŞ ýùÀŁĂŁĂŁîŘÀŁŻ’ ø (U+2028) "ŚŁûÂŁÎŁĺ ýùÀŁĂŁĂŁîŘÀŁŻ’ ýŚŁûĆŁÆŁþĄŁœ ěŘ ÀŁþŚŁśŁœ ,ÀŁĂŁî ÀŁČŁóĄŁŮ UTF-8 řc ŁóŚŁì ÀŁĂŁî ùćŚŁÔŁţŁĺŘ ćÂŁŞďŚŁî Â cŁţŁÆŁŞ ďć ŚŁûÂŁÎŁĺ ö cćÂŁîŘÀŁŻ ń c ĄŁÊŁ£Łõ ťc ŁõĚŁä ěŘ ÀŁþŚŁŞ ĆŁØŁÜŁŞ .ÀŁĂŁî ùćŚŁÔŁţŁĺŘ (U+2029)

.ťŁĺŘ ⟨U+000D, U+000A⟩ ŚŁþ ,U+000D ,U+000A fĎĄŁÞŁãŁõ ĆŁî

Ÿ ťc ĺĄČŢ

ýěďŘÝŁû ø ýěŚŁĺñŚŁõÂŁœ

(üŁŮŚŁäĚŁÏŘ)

ÛŁŁþÀŁŁśŁŁŮ ŚŁŁûĆŁŁÆŁŁþĄŁŁœ ěŘ ýŘĆŁŁţŁŁłď ĆŁŁŞ üŁŁÔŁŁÜŁŁţŁŁ£ŁŁõ ýŚŁŁûľøď ĆŁŁŞ ÀŁŁœŘĄŁŁŮüŁŁõ ŚŁŁûćŚŁŁÞŁŁœ ø éøÂŁŁ§ ěŘ ýŘĆŁŁóŚŁŁśŁŁœć ĆŁŁî ŚŁŁźŁŁœŊ ěŘ ýŚŁûĆŁÆŁþĄŁœ ýĆŁóŚŁśŁœć ŚŁŞ ÝŁû ø ,"ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ øŘø’ ýĆŁÆŁþĄŁœ ŚŁŞ ÝŁû öŘĄŁŮüŁõ Řď "SSŁõĄbŁõ’ ýĆŁÞŁÜŁî fĚŁůŁõ) ćĄŁł Řď ŚŁûöŊ ,ćĄŁą ýŚŁûüŁŻøÂŁą ü c ŁÚŁĂŁûŚŁÞŁû ď cĄŁÑŁĂŁõĆŁŞ ťŁĺŘ ÂŁţŁúŁŞ ŚŁûćÂŁŞďŚŁî ,(ćÂŁî ýďŘÁŁðÀŁîe "ĎŚŁŞ ýùÃŁÞŁû ,øŘø’ ťŁĺŘ ùÀŁł ÓŁČŁňĄŁŮ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ 15 ýĆŁÞŁČŁÞŁő ďć ĆŁî ,ÀŁîüŁœĄŁþ ñ cŚŁõÂŁœ ýŚŁûŰďĄŁň ěŘ üŁØŁþ Ű cďĄŁňĆŁŞ

.ÀŁĂŁĂŁî ÀŁČŁóĄŁŮ

(Normalizaion Form C) C ñ cŚŁŁõÂŁŁœ Ű cďĄŁŁň ,ŚŁŁûùćŘć ñ cćŚŁŁśŁŁŮ ďć ĳŁŁŻaÂŁŁõ řc ŁŁóŚŁŁì ö cŘĄŁŁĂŁŁäĆŁŁŞ ,ÂŁŁőŚŁŁ§ ćcďŘÀŁŁœŚŁŁţŁŁĺŘ ďć öŊ ö cÀŁŁłĆŁŁČŁŁňĄŁŁŮ ø ,řŁŁóŚŁŁì SSŁŁþŘ ö cćĄŁŁŞ ÂŁŁČŁŁðŘÂŁŁê ťc ŁŁÜŁŁä ĆŁŁŞ ŠŚŁŁ£ŁŁţŁŁœŘ SSŁŁþŘ .ÀŁŁĂŁŁîüŁŁõ ĆŁŁČŁŁňĄŁŁŮ Řď UTF-8 řc ŁŁóŚŁŁì ø

.ťŁĺŘ ĆŁţŁêÂŁð ŰďĄŁň ,XML ø HTML ĆŁÜŁÞŁŻ ěŘ ,World Wide Web ô cĄŁČŁĺÂŁÆŁĂŁî ýŚŁûćďŘÀŁœŚŁţŁĺŘ Û cŁŁŁØŁŁŁł ĆŁŁŁŞ ,ÀŁŁŁœĄŁŁŁł ýďŘÁŁŁŁðÀŁŁŁî ÓŁŁŁÜŁŁŁţŁŁŁ£ŁŁŁõ Û cŁŁŁØŁŁŁł ÀŁŁŁĂŁŁŁİ ĆŁŁŁŞ ÀŁŁŁĂŁŁŁœŘĄŁŁŁŮüŁŁŁõ ĆŁŁŁî üŁŁŁþŚŁŁŁûĆŁŁŁÆŁŁŁþĄŁŁŁœ ,"C ñ cŚŁŁŁõÂŁŁŁœ Ű cďĄŁŁŁň’ ďć ťc ŁŁŁŁŁõĚŁŁŁŁŁä ÀŁŁŁŁŁĂŁŁŁŁŁİ ĆŁŁŁŁŁî üŁŁŁŁŁþŚŁŁŁŁŁûťŁŁŁŁŁóŚŁŁŁŁŁ§ ýŘÂŁŁŁŁŁŞ ýÀŁŁŁŁŁ§Řø řc ŁŁŁŁŁČŁŁŁŁŁŮÂŁŁŁŁŁŮ ,ùøĚŁŁŁŁŁäĆŁŁŁŁŁŞ .ÀŁŁŁŁŁœĄŁŁŁŁŁłüŁŁŁŁŁõ ýďŘÁŁŁŁŁŁðÀŁŁŁŁŁî ýŘĆŁŁŁŁŁÆŁŁŁŁŁþĄŁŁŁŁŁœŒŁŁŁŁŁŮ

.ćĄŁłüŁõ SSŁČŁČŁãŁŮ ÀŁœÂŁČŁðüŁõ ďŘÂŁì üŁĺÂŁî éc ÂŁ§ ŒŁþ ýøď ùÀŁœĄŁłřŁČŁîÂŁŮ

ťŁŁĺŘ ÂŁŁţŁŁúŁŁŞ ,ÀŁŁĂŁŁĂŁŁîüŁŁõ üŁŁœŚŁŁśŁŁČŁŁţŁŁÈŁŁŢ Řď üŁŁÔŁŁÜŁŁţŁŁ£ŁŁõ ñ cŚŁŁõÂŁŁœ ýŚŁŁûŰďĄŁŁň ŚŁŁûćÂŁŁŞďŚŁŁî ĆŁŁî üŁŁŮďĄŁŁň ďć ,ÂŁŁÚŁŁþć éc ÂŁŁÏ ěŘ ÀŁŁŁîüŁŁŁœĄŁŁŁþ ćcďŘÀŁŁŁœŚŁŁŁţŁŁŁĺŘ 3 ø 2 ýŚŁŁŁûÛŁŁŁÊŁŁŁê ĆŁŁŁŞ ,ÂŁŁŁţŁŁŁÈŁŁŁČŁŁŁŞ á c ĚŁŁŁÏŘ ýŘÂŁŁŁŞ .ÀŁŁŁĂŁŁŁûć ÉŁŁŁČŁŁŁ£ŁŁŁÈŁŁŁŮ Řď "ěďŘÝŁŁŁû’ ýŚŁŁŁûĆŁŁŁţŁŁŁłď

.ÀŁČŁĂŁî ĆŁãŁŻŘÂŁõ

fĚŁŁŁůŁŁŁõ ,ÀŁŁŁĂŁŁŁĂŁŁŁî üŁŁŁœŚŁŁŁśŁŁŁČŁŁŁţŁŁŁÈŁŁŁŢ ÃŁŁŁČŁŁŁœ Řď "ýěďŘÝŁŁŁû’ ěŘ ýÂŁŁŁŮĎŚŁŁŁŞ ýŚŁŁŁûĆŁŁŁþĎ ÀŁŁŁĂŁŁŁœŘĄŁŁŁŮüŁŁŁõ ŚŁŁŁûćÂŁŁŁŞďŚŁŁŁî ŃŃŃ ýďøŊćŚŁŁŁþ ýùÃŁŁÞŁŁû ,ý c ’ ýŚŁŁûĆŁŁÆŁŁþĄŁŁœĆŁŁţŁŁłď SS cŁŁČŁŁŞ ŚŁŁþ ,"üŁŁŞÂŁŁä éc ŚŁŁî’ ø "éŚŁŁî’ ÀŁŁĂŁŁœŚŁŁõ üŁŁþŚŁŁûĆŁŁÆŁŁþĄŁŁœ SS cŁŁČŁŁŞ ÓŁŁČŁŁãŁŁő ý c ěďŘÝŁŁû

.ćěŘćÂŁŢüŁÞŁœ ŚŁûýěďŘÝŁû ĆŁœĄŁð SSŁþŘ ĆŁŞ ćďŘÀŁœŚŁţŁĺŘ SSŁþŘ ."ĎŚŁŞ ýùÃŁÞŁû ,ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c ’ ø "ĎŚŁŞ

Ž ťc ĺĄČŢ

ĆŁõŚŁœùęŘø

(üŁŮŚŁäĚŁÏŘ)

right-joining . . . . . . . . . . . . . . . . . . . . . . . . ÛŁňøťŁĺŘď string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁţŁłď embedding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . SSŁţŁõÂŁþě subtext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . SSŁţŁõÂŁþě data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ùćŘć insert. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Žďć device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ùŚŁÚŁţŁĺć dual-joining . . . . . . . . . . . . . . . . . . . . . . . . . . . ÛŁňøøć storage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . ýěŚŁĺùÂŁČŁąč conformant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ďŚŁðěŚŁĺ line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÂŁÎŁĺ line-breaking . . . . . . . . . . . . . . . . . . . . . . . . üŁĂŁØŁłÂŁÎŁĺ transparent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . éŚŁÔŁł glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÛŁØŁł final form. . . . . . . . . . . . . . . . . . . . . . . . . . . . ÂŁąŊ Û cŁØŁł initial form. . . . . . . . . . . . . . . . . . . . . . . . . . . ñøŘ Û cŁØŁł isolated form. . . . . . . . . . . . . . . . . . . . . . . . . ŚŁúŁĂŁŮ Û cŁØŁł presentation form . . . . . . . . . . . . . . . . . üŁÈŁþŚŁÞŁœ Û cŁØŁł medial form . . . . . . . . . . . . . . . . . . . . . . . . ÍŁĺø Û cŁØŁł Basic Multilingual Plane . . . . . . . . . . ĆŁþŚŁŢ ýĆŁ¡ŁÔŁň sign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ťŁõĚŁä presentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁŐŘďŘ information. . . . . . . . . . . . . . . . . . . . . . . . . . . ŰŚŁäĚŁÏŘ Cursive Joining Algorithm . . . . . . ñŚŁÊŁŮŘ Ý cŁţŁþďĄŁÚŁóŘ Bidirectional Algorithm . . . . . . . . ĆŁţŁúŁŻøć Ý cŁţŁþďĄŁÚŁóŘ transfer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ñŚŁÖŁţŁœŘ ideographic . . . . . . . . . . . . . . . . . . . . . . . . . ďŚŁÚŁœĆŁÈŁþÀŁœŘ octet/byte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ťŁþŚŁŞ platform. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÂŁţŁÆŁŞ block . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . íĄŁÜŁŞ paragraph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÀŁĂŁŞ processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ľěŘćÂŁŢ file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ùÀŁœøÂŁŢ support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁœŚŁśŁČŁţŁÈŁŢ conform . . . . . . . . . . . . . . . . . . . . . . . . . . öćÂŁî ýøďüŁŢ interchange . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ñćŚŁśŁŮ visual order . . . . . . . . . . . . . . . . . . . . . . ýďŘÀŁþć řc ŁČŁŮÂŁŮ logical order . . . . . . . . . . . . . . . . . . . . . . üŁþŚŁĂŁãŁõ řc ŁČŁŮÂŁŮ left-joining. . . . . . . . . . . . . . . . . . . . . . . . . . . ÛŁňøšŁİ delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . éÁŁ§ letter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . éÂŁ§ script . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÍŁą character set . . . . . . . . . . . . . . . . . . . . . ĆŁÆŁþĄŁœĆŁäĄŁÞŁźŁõ environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÍŁČŁ¡Łõ sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ýěŚŁĺřŁŮÂŁõ symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ćŚŁÞŁœ display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÇŁþŚŁÞŁœ document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ďŚŁţŁłĄŁœ . . . . . . . . . . . . . . . . . . . ŋÂŁäüŁŞ/ùÀŁœĄŁłřŁČŁîÂŁŮ ýĆŁÆŁþĄŁœ combining/non-spacing character . . . . . . . . . . . . . private use character . . . . . . . . . üŁňĄŁÊŁą ýĆŁÆŁþĄŁœ base letter/character. . . . . . . . . . . . . üŁĺÂŁî ýĆŁÆŁþĄŁœ control character . . . . . . . . . . . . . . . . üŁóÂŁţŁĂŁî ýĆŁÆŁþĄŁœ entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ýěŚŁĺćďŘø join-causing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÛŁňŘø edit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÇŁþŘÂŁþø non-joining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÛŁňŚŁê compression. . . . . . . . . . . . . . . . . . . . . . . . ýěŚŁĺùćÂŁÈŁê encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . řŁóŚŁì transformation format . . . . . . . . . . . . . . ñćŚŁśŁŮ řc ŁóŚŁì mirroring . . . . . . . . . . . . . . . . . . . . . . . . . . . ýěŚŁĺĆŁĂŁþÂŁì font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÝŁÜŁì application . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ćÂŁŞďŚŁî code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÀŁîe to encode . . . . . . . . . . . . . . . . . . . . . . . öćÂŁî ýďŘÁŁðÀŁîe ligature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ďĄŁŮŚŁÚŁČŁó text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . SSŁţŁõ plain text . . . . . . . . . . . . . . . . . . . . . . . . . . . . ùćŚŁĺ SS cŁţŁõ left join-causing . . . . . . . . . . . . . . . . . . . šŁİĆŁŞÛŁÊŁţŁõ right join-causing . . . . . . . . . . . . . . . . ťŁĺŘďĆŁŞÛŁÊŁţŁõ

Ĳ ťc ĺĄČŢ

ŚŁûĆŁÆŁþĄŁœ À cŁî

(üŁŮŚŁäĚŁÏŘ)

ü c ŁþŚŁśŁÔŁóŘ řc ŁČŁŮÂŁŮĆŁŞ ĆŁî ťŁĺŘ ÀŁîüŁœĄŁþ ćcďŘÀŁœŚŁţŁĺŘ ďć ,ÂŁőŚŁ§ ćcďŘÀŁœŚŁţŁĺŘ ýŚŁûĆŁÆŁþĄŁœ ýŚŁûÀŁî ťc ŁĺÂŁúŁê ,ťŁĺĄŁČŁŢ SSŁþŘ .ťŁĺŘ ùÀŁł řŁŮÂŁõ ĆŁÆŁþĄŁœ ô cŚŁœ

0623. . . . . . . . . . . . . . .ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ ü c ŁĺďŚŁê éc ÂŁ§ 0628 . . . . . . . . . . . . . . . . . . . . . . . . . . . . Š c ü c ŁĺďŚŁê éc ÂŁ§ 067E . . . . . . . . . . . . . . . . . . . . . . . . . . . . Ť c ü c ŁĺďŚŁê éc ÂŁ§ 062A . . . . . . . . . . . . . . . . . . . . . . . . . . . . Ű c ü c ŁĺďŚŁê éc ÂŁ§ 062B . . . . . . . . . . . . . . . . . . . . . . . . . . . . Ÿ c ü c ŁĺďŚŁê éc ÂŁ§ 062C . . . . . . . . . . . . . . . . . . . . . . . . . . . ÝŁČŁŻ ü c ŁĺďŚŁê éc ÂŁ§ 0686. . . . . . . . . . . . . . . . . . . . . . . . . . . . .Ĳ c ü c ŁĺďŚŁê éc ÂŁ§ 062D. . . . . . . . . . . . . . . . . . . . . . . . . . . . .đ c ü c ŁĺďŚŁê éc ÂŁ§ 062E. . . . . . . . . . . . . . . . . . . . . . . . . . . . .ă c ü c ŁĺďŚŁê éc ÂŁ§ 062F . . . . . . . . . . . . . . . . . . . . . . . . . . . ñŘć ü c ŁĺďŚŁê éc ÂŁ§ 0630 . . . . . . . . . . . . . . . . . . . . . . . . . . . ñŘč ü c ŁĺďŚŁê éc ÂŁ§ 0631 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ď cü c ŁĺďŚŁê éc ÂŁ§ 0632 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ěc ü c ŁĺďŚŁê éc ÂŁ§ 0698 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ęc ü c ŁĺďŚŁê éc ÂŁ§ 0633 . . . . . . . . . . . . . . . . . . . . . . . . . . SSŁČŁĺ ü c ŁĺďŚŁê éc ÂŁ§ 0634 . . . . . . . . . . . . . . . . . . . . . . . . . . SSŁČŁł ü c ŁĺďŚŁê éc ÂŁ§ 0635. . . . . . . . . . . . . . . . . . . . . . . . . . .ćŚŁň ü c ŁĺďŚŁê éc ÂŁ§ 0636. . . . . . . . . . . . . . . . . . . . . . . . . . .ćŚŁő ü c ŁĺďŚŁê éc ÂŁ§ 007B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ěŚŁŞ ćĎĄŁî Ŋ 007D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁţŁÆŁŞ ćĎĄŁî Ŋ 200D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ýěŚŁźŁõ ñ cŚŁÊŁŮŘ 0670. . . . . . . . . . . . . . . . . . . . . . .üŁĺďŚŁê ýùďĄŁÊŁÖŁõ Ó c ŁóŘ 202C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . SSŁţŁõÂŁþě ö cŚŁþŚŁŢ 0028. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .ěŚŁŞ ÃŁţŁœŘÂŁŢ 0029 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁţŁÆŁŞ ÃŁţŁœŘÂŁŢ 064F. . . . . . . . . . . . . . . . . . . . . . . . .(ĆŁÞŁő) üŁĺďŚŁê Ç c ŁČŁŢ 0651 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê À cŁþÀŁÈŁŮ 002D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁúŁĂŁõùÂŁČŁŮ 2029. . . . . . . . . . . . . . . . . . . . . . . . . . .ŚŁûÀŁĂŁŞ ýùÀŁĂŁĂŁîŘÀŁŻ 2028 . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁûÂŁÎŁĺ ýùÀŁĂŁĂŁîŘÀŁŻ 066C . . . . . . . . . . . . . . . . . üŁĺďŚŁê ýŚŁûďŘÃŁû ýùÀŁĂŁĂŁîŘÀŁŻ 0625 . . . . . . . . . . . . . . . . . . . SSŁČŁþŚŁŢ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ éc ÂŁ§ 0671. . . . . . . . . . . . . . . . . . . . . . . . . . . .ÛŁňø Ó c ŁóŘ éc ÂŁ§ 0629. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .ćÂŁð Ű c éc ÂŁ§ 0622 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Ŋ ü c ŁĺďŚŁê éc ÂŁ§ 0627 . . . . . . . . . . . . . . . . . . . . . . . . . . . ÓŁóŘ ü c ŁĺďŚŁê éc ÂŁ§ 064D . . . . . . . . . . . . . . . . . . . . (ÂŁŻ SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Â cŁþěøć 003A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁÎŁÖŁœøć 06F5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÿŁĂŁŢ ü c ŁĺďŚŁê Ý cŁìď 06F4. . . . . . . . . . . . . . . . . . . . . . . . . . . .ďŚŁúŁİ ü c ŁĺďŚŁê Ý cŁìď 06F2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . øć ü c ŁĺďŚŁê Ý cŁìď 06F3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁĺ ü c ŁĺďŚŁê Ý cŁìď 06F6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÇŁł ü c ŁĺďŚŁê Ý cŁìď 06F0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÂŁÔŁň ü c ŁĺďŚŁê Ý cŁìď 06F9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁœ ü c ŁĺďŚŁê Ý cŁìď 06F8 . . . . . . . . . . . . . . . . . . . . . . . . . . . ťŁÈŁû ü c ŁĺďŚŁê Ý cŁìď 06F7. . . . . . . . . . . . . . . . . . . . . . . . . . . .ťŁÔŁû ü c ŁĺďŚŁê Ý cŁìď 06F1. . . . . . . . . . . . . . . . . . . . . . . . . . . . .ŒŁþ ü c ŁĺďŚŁê Ý cŁìď 064E . . . . . . . . . . . . . . . . . . . . . . . . . . (ĆŁ¡ŁţŁê) üŁĺďŚŁê Â cŁŞdě 0650 . . . . . . . . . . . . . . . . . . . . . . . . . . (ùÂŁÆŁî) üŁĺďŚŁê Â cŁþě 202D . . . . . . . . . . . . . . . . . . . . ťŁĺŘďĆŁŞšŁİ fŘÀŁČŁîŘ SS cŁţŁõÂŁþě 202E . . . . . . . . . . . . . . . . . . . . šŁİĆŁŞťŁĺŘď fŘÀŁČŁîŘ SS cŁţŁõÂŁþě 202A. . . . . . . . . . . . . . . . . . . . . . . . .ťŁĺŘďĆŁŞšŁİ SS cŁţŁõÂŁþě 202B. . . . . . . . . . . . . . . . . . . . . . . . .šŁİĆŁŞťŁĺŘď SS cŁţŁõÂŁþě 0652. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .üŁĺďŚŁê SS cŁîŚŁĺ 002A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ùďŚŁţŁĺ 000D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÂŁÎŁĺ Â cŁĺ 000A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÀŁãŁŞ Â cŁÎŁĺ 2026 . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁÖŁêŘ ýĆŁÎŁÖŁœĆŁĺ 0020 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁÜŁňŚŁê 200C . . . . . . . . . . . . . . . . . . . . . . . . . . . ýěŚŁźŁõ ýĆŁÜŁňŚŁê 005B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ěŚŁŞ ĆŁłøÂŁî 005D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁţŁÆŁŞ ĆŁłøÂŁî 0637 . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁÏ ü c ŁĺďŚŁê éc ÂŁ§ 0638 . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁÒ ü c ŁĺďŚŁê éc ÂŁ§ 0639. . . . . . . . . . . . . . . . . . . . . . . . . . .SSŁČŁä ü c ŁĺďŚŁê éc ÂŁ§ 063A. . . . . . . . . . . . . . . . . . . . . . . . . . .SSŁČŁè ü c ŁĺďŚŁê éc ÂŁ§ 0641 . . . . . . . . . . . . . . . . . . . . . . . . . . . . éc ü c ŁĺďŚŁê éc ÂŁ§ 0642. . . . . . . . . . . . . . . . . . . . . . . . . . .éŚŁì ü c ŁĺďŚŁê éc ÂŁ§ 06A9 . . . . . . . . . . . . . . . . . . . . . . . . . . éŚŁî ü c ŁĺďŚŁê éc ÂŁ§ 06AF . . . . . . . . . . . . . . . . . . . . . . . . . . éŚŁð ü c ŁĺďŚŁê éc ÂŁ§ 0644. . . . . . . . . . . . . . . . . . . . . . . . . . . .ôĎ ü c ŁĺďŚŁê éc ÂŁ§ 0645 . . . . . . . . . . . . . . . . . . . . . . . . . . . ÝŁČŁõ ü c ŁĺďŚŁê éc ÂŁ§ 0646 . . . . . . . . . . . . . . . . . . . . . . . . . . . öĄŁœ ü c ŁĺďŚŁê éc ÂŁ§ 0648 . . . . . . . . . . . . . . . . . . . . . . . . . . . . øŘø ü c ŁĺďŚŁê éc ÂŁ§ 0624. . . . . . . . . . . . . . . .ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ øŘø ü c ŁĺďŚŁê éc ÂŁ§ 0647 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Łû c ü c ŁĺďŚŁê éc ÂŁ§ 0621 . . . . . . . . . . . . . . . . . . . . . . . . . . ùÃŁÞŁû ü c ŁĺďŚŁê éc ÂŁ§ 06CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . ý c ü c ŁĺďŚŁê éc ÂŁ§ 0626 . . . . . . . . . . . . . . . . ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ý c ü c ŁĺďŚŁê éc ÂŁ§ 0643 . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁŞÂŁä éc ŚŁî éc ÂŁ§ 0649. . . . . . . . . . . . . . . . . . . . . .ĆŁÎŁÖŁœüŁŞ ü c ŁŞÂŁä ý c éc ÂŁ§ 064A. . . . . . . . . . . . . . . . . . . . . .ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c éc ÂŁ§ 002F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . řŁþďŘ Íc Łą 005C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . øďŘø řc ŁþďŘ Íc Łą 2010 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ùÂŁČŁŮ Íc Łą 007C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ýćĄŁÞŁä Íc Łą 066A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê À cŁňďć 064C . . . . . . . . . . . . . . . . . . (âŁêď SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Ç c ŁČŁŢøć 064B . . . . . . . . . . . . . . . . . . (řŁÊŁœ SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Â cŁŞděøć 2212 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŚŁúŁĂŁõ ťc ŁõĚŁä 0653 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê À cŁõ 066B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê Ã cŁČŁÞŁõ FEFF . . . . . . . . . . . . . . . . . . . . . . . ŚŁûťŁþŚŁŞ řc ŁČŁŮÂŁŮ ýĆŁœŚŁÈŁœ 200E . . . . . . . . . . . . . . . . . . . . . . . . ťŁĺŘďĆŁŞšŁİ ýĆŁœŚŁÈŁœ 200F . . . . . . . . . . . . . . . . . . . . . . . . šŁİĆŁŞťŁĺŘď ýĆŁœŚŁÈŁœ 002E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁÎŁÖŁœ 061B . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ñ cĄŁðÂŁþøĆŁÎŁÖŁœ 060C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ñ cĄŁðÂŁþø 0654 . . . . . . . . . . . . . . . . . . . . . . . . . . ĎŚŁŞ ü c ŁĺďŚŁê ýùÃŁÞŁû 0655 . . . . . . . . . . . . . . . . . . . . . . . . SSŁČŁþŚŁŢ ü c ŁĺďŚŁê ýùÃŁÞŁû 0640 . . . . . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ü c ŁðÀŁČŁÈŁî 00AB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ěŚŁŞ ĆŁõĄŁČŁð 00BB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁţŁÆŁŞ ĆŁõĄŁČŁð 003E. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .ÂŁţŁðďÃŁŞ ťc ŁõĚŁä 002B . . . . . . . . . . . . . . . . . . . . . . . . . . . . ĆŁêŚŁőŘĆŁŞ ťc ŁõĚŁä 0021 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . řŁźŁãŁŮ ťc ŁõĚŁä 00F7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÝŁČŁÆŁÖŁŮ ťc ŁõĚŁä 061F . . . . . . . . . . . . . . . . . . . . . . . . üŁĺďŚŁê ñ cŘĄbŁĺ ťc ŁõĚŁä 00D7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ŠÂŁő ťc ŁõĚŁä 003C . . . . . . . . . . . . . . . . . . . . . . . . . . . . ÂŁţŁØŁİĄŁî ťc ŁõĚŁä 003D. . . . . . . . . . . . . . . . . . . . . . . . . . . . .ýøŚŁÆŁõ ťc ŁõĚŁä

đ ťc ĺĄČŢ

ŚŁûĆŁÆŁþĄŁœ ô cŚŁœ

(üŁŮŚŁäĚŁÏŘ)

ćcďŘÀŁœŚŁţŁĺŘ ďć ĆŁÆŁþĄŁœ À cŁî řc ŁČŁŮÂŁŮĆŁŞ ,ÂŁőŚŁ§ ćcďŘÀŁœŚŁţŁĺŘ ďć ùÀŁłÓŁþÂŁãŁŮ ýŚŁûĆŁÆŁþĄŁœ ýŚŁûôŚŁœ ťc ŁĺÂŁúŁê ,ťŁĺĄŁČŁŢ SSŁþŘ

.ťŁĺŘ ÀŁîüŁœĄŁþ

ěŚŁŞ ćĎĄŁî Ŋ 007B ýćĄŁÞŁä Íc Łą 007C ĆŁţŁÆŁŞ ćĎĄŁî Ŋ 007D

ěŚŁŞ ĆŁõĄŁČŁð 00AB ĆŁţŁÆŁŞ ĆŁõĄŁČŁð 00BB

ŠÂŁő ťc ŁõĚŁä 00D7 ÝŁČŁÆŁÖŁŮ ťc ŁõĚŁä 00F7* üŁĺďŚŁê ñ cĄŁðÂŁþø 060C

üŁĺďŚŁê ñ cĄŁðÂŁþøĆŁÎŁÖŁœ 061B üŁĺďŚŁê ñ cŘĄbŁĺ ťc ŁõĚŁä 061F ùÃŁÞŁû ü c ŁĺďŚŁê éc ÂŁ§ 0621

Ŋ ü c ŁĺďŚŁê éc ÂŁ§ 0622

ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ ü c ŁĺďŚŁê éc ÂŁ§ 0623 ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ øŘø ü c ŁĺďŚŁê éc ÂŁ§ 0624

SSŁČŁþŚŁŢ ýùÃŁÞŁû ŚŁŞ ÓŁóŘ éc ÂŁ§ 0625* ĎŚŁŞ ýùÃŁÞŁû ŚŁŞ ý c ü c ŁĺďŚŁê éc ÂŁ§ 0626

ÓŁóŘ ü c ŁĺďŚŁê éc ÂŁ§ 0627 Š c ü c ŁĺďŚŁê éc ÂŁ§ 0628 ÀŁãŁŞ Â cŁÎŁĺ 000A ÂŁÎŁĺ Â cŁĺ 000D ĆŁÜŁňŚŁê 0020

řŁźŁãŁŮ ťc ŁõĚŁä 0021 ěŚŁŞ ÃŁţŁœŘÂŁŢ 0028 ĆŁţŁÆŁŞ ÃŁţŁœŘÂŁŢ 0029 ùďŚŁţŁĺ 002A

ĆŁêŚŁőŘĆŁŞ ťc ŁõĚŁä 002B

ŚŁúŁĂŁõùÂŁČŁŮ 002D ĆŁÎŁÖŁœ 002E řŁþďŘ Íc Łą 002F ĆŁÎŁÖŁœøć 003A

ÂŁţŁØŁİĄŁî ťc ŁõĚŁä 003C ýøŚŁÆŁõ ťc ŁõĚŁä 003D ÂŁţŁðďÃŁŞ ťc ŁõĚŁä 003E ěŚŁŞ ĆŁłøÂŁî 005B øďŘø řc ŁþďŘ Íc Łą 005C ĆŁţŁÆŁŞ ĆŁłøÂŁî 005D ĆŁÎŁÖŁœüŁŞ ü c ŁŞÂŁä ý c éc ÂŁ§ 0649* ďŘćĆŁÎŁÖŁœ ü c ŁŞÂŁä ý c éc ÂŁ§ 064A*

(řŁÊŁœ SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Â cŁŞděøć 064B (âŁêď SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Ç c ŁČŁŢøć 064C

(ÂŁŻ SS cŁþĄŁĂŁŮ) üŁĺďŚŁê Â cŁþěøć 064D (ĆŁ¡ŁţŁê) üŁĺďŚŁê Â cŁŞdě 064E (ĆŁÞŁő) üŁĺďŚŁê Ç c ŁČŁŢ 064F (ùÂŁÆŁî) üŁĺďŚŁê Â cŁþě 0650

üŁĺďŚŁê À cŁþÀŁÈŁŮ 0651 üŁĺďŚŁê SS cŁîŚŁĺ 0652

üŁĺďŚŁê À cŁõ 0653* ĎŚŁŞ ü c ŁĺďŚŁê ýùÃŁÞŁû 0654 SSŁČŁþŚŁŢ ü c ŁĺďŚŁê ýùÃŁÞŁû 0655*

üŁĺďŚŁê À cŁňďć 066A üŁĺďŚŁê Ã cŁČŁÞŁõ 066B

üŁĺďŚŁê ýŚŁûďŘÃŁû ýùÀŁĂŁĂŁîŘÀŁŻ 066C

üŁĺďŚŁê ýùďĄŁÊŁÖŁõ Ó c ŁóŘ 0670 ÛŁňø Ó c ŁóŘ éc ÂŁ§ 0671* Ť c ü c ŁĺďŚŁê éc ÂŁ§ 067E Ĳ c ü c ŁĺďŚŁê éc ÂŁ§ 0686 ęc ü c ŁĺďŚŁê éc ÂŁ§ 0698

éŚŁî ü c ŁĺďŚŁê éc ÂŁ§ 06A9 éŚŁð ü c ŁĺďŚŁê éc ÂŁ§ 06AF ý c ü c ŁĺďŚŁê éc ÂŁ§ 06CC ÂŁÔŁň ü c ŁĺďŚŁê Ý cŁìď 06F0 ŒŁþ ü c ŁĺďŚŁê Ý cŁìď 06F1 øć ü c ŁĺďŚŁê Ý cŁìď 06F2

ćÂŁð Ű c éc ÂŁ§ 0629 Ű c ü c ŁĺďŚŁê éc ÂŁ§ 062A Ÿ c ü c ŁĺďŚŁê éc ÂŁ§ 062B ÝŁČŁŻ ü c ŁĺďŚŁê éc ÂŁ§ 062C

đ c ü c ŁĺďŚŁê éc ÂŁ§ 062D ă c ü c ŁĺďŚŁê éc ÂŁ§ 062E ñŘć ü c ŁĺďŚŁê éc ÂŁ§ 062F ñŘč ü c ŁĺďŚŁê éc ÂŁ§ 0630

ď cü c ŁĺďŚŁê éc ÂŁ§ 0631 ěc ü c ŁĺďŚŁê éc ÂŁ§ 0632

SSŁČŁĺ ü c ŁĺďŚŁê éc ÂŁ§ 0633 SSŁČŁł ü c ŁĺďŚŁê éc ÂŁ§ 0634 ćŚŁň ü c ŁĺďŚŁê éc ÂŁ§ 0635 ćŚŁő ü c ŁĺďŚŁê éc ÂŁ§ 0636 ŚŁÏ ü c ŁĺďŚŁê éc ÂŁ§ 0637 ŚŁÒ ü c ŁĺďŚŁê éc ÂŁ§ 0638 SSŁČŁä ü c ŁĺďŚŁê éc ÂŁ§ 0639 SSŁČŁè ü c ŁĺďŚŁê éc ÂŁ§ 063A üŁĺďŚŁê ü c ŁðÀŁČŁÈŁî 0640 éc ü c ŁĺďŚŁê éc ÂŁ§ 0641 éŚŁì ü c ŁĺďŚŁê éc ÂŁ§ 0642 üŁŞÂŁä éc ŚŁî éc ÂŁ§ 0643* ôĎ ü c ŁĺďŚŁê éc ÂŁ§ 0644 ÝŁČŁõ ü c ŁĺďŚŁê éc ÂŁ§ 0645 öĄŁœ ü c ŁĺďŚŁê éc ÂŁ§ 0646

Łû c ü c ŁĺďŚŁê éc ÂŁ§ 0647 øŘø ü c ŁĺďŚŁê éc ÂŁ§ 0648

ùÂŁČŁŮ Íc Łą 2010* üŁÖŁêŘ ýĆŁÎŁÖŁœĆŁĺ 2026*

ŚŁûÂŁÎŁĺ ýùÀŁĂŁĂŁîŘÀŁŻ 2028* ŚŁûÀŁĂŁŞ ýùÀŁĂŁĂŁîŘÀŁŻ 2029*

ťŁĺŘďĆŁŞšŁİ SS cŁţŁõÂŁþě 202A* šŁİĆŁŞťŁĺŘď SS cŁţŁõÂŁþě 202B*

SSŁţŁõÂŁþě ö cŚŁþŚŁŢ 202C*

ťŁĺŘďĆŁŞšŁİ fŘÀŁČŁîŘ SS cŁţŁõÂŁþě 202D* šŁİĆŁŞťŁĺŘď fŘÀŁČŁîŘ SS cŁţŁõÂŁþě 202E*

ŚŁúŁĂŁõ ťc ŁõĚŁä 2212* ŚŁûťŁþŚŁŞ řc ŁČŁŮÂŁŮ ýĆŁœŚŁÈŁœ FEFF ĆŁĺ ü c ŁĺďŚŁê Ý cŁìď 06F3 ďŚŁúŁİ ü c ŁĺďŚŁê Ý cŁìď 06F4 ÿŁĂŁŢ ü c ŁĺďŚŁê Ý cŁìď 06F5 ÇŁł ü c ŁĺďŚŁê Ý cŁìď 06F6 ťŁÔŁû ü c ŁĺďŚŁê Ý cŁìď 06F7 ťŁÈŁû ü c ŁĺďŚŁê Ý cŁìď 06F8

ĆŁœ ü c ŁĺďŚŁê Ý cŁìď 06F9 ýěŚŁźŁõ ýĆŁÜŁňŚŁê 200C ýěŚŁźŁõ ñ cŚŁÊŁŮŘ 200D

ťŁĺŘďĆŁŞšŁİ ýĆŁœŚŁÈŁœ 200E šŁİĆŁŞťŁĺŘď ýĆŁœŚŁÈŁœ 200F

.ťŁĺŘ ĆŁÆŁþĄŁœ öŊ ö cćĄŁŞ ýďŚŁČŁţŁąŘ ýŚŁĂŁãŁõĆŁŞ ĆŁÆŁþĄŁœ À cŁî ťc ŁĺŘď ťc ŁÞŁĺ ďć ùďŚŁţŁĺ ťc ŁõĚŁä ŃŃŃ ýďøŊćŚŁþ